### v1

# Universal Approximation Theory: Foundations for Parallelism in Neural Networks

Abstract
Neural networks are increasingly evolving towards training large models with big data, a method that has demonstrated superior performance across many tasks. However, this approach introduces an urgent problem: current deep learning models are predominantly serial, meaning that as the number of network layers increases, so do the training and inference times. This is unacceptable if deep learning is to continue advancing. Therefore, this paper proposes a deep learning parallelization strategy based on the Universal Approximation Theorem (UAT). From this foundation, we designed a parallel network called Para-Former to test our theory. Unlike traditional serial models, the inference time of Para-Former does not increase with the number of layers, significantly accelerating the inference speed of multi-layer networks. Experimental results validate the effectiveness of this network.
1 Introduction
In today’s era, deep learning models have undoubtedly taken the lead in the fields of computer vision (CV) and natural language processing (NLP), demonstrating exceptional performance across numerous tasks. In CV, models such as ResNet(He et al. 2015) and ViT(Vaswani et al. 2017) for image recognition, U-Net (Ronneberger, Fischer, and Brox 2015) for image segmentation, GAN (Goodfellow et al. 2014) for image generation, and diffusion models like SORA (Liu et al. 2024) for video generation highlight the extensive application of deep learning in CV. These models are primarily based on convolutional neural networks (CNNs), Transformers, or a combination of both. Simultaneously, large language models (LLMs) based on Transformers have become dominant in natural language processing. Overall, deep learning models have shown vast application potential and powerful capabilities so far.
However, this progress has brought forth some challenges: the development trend in deep learning is moving towards larger models and larger datasets. This trend is fundamentally driven by the UAT(Cybenko 2007; Hornik, Stinchcombe, and White 1989). From the papers UAT2LLM(Wang and Li 2024) and UAT2CV(Wang and Li 2024), we know that residual-based CNNs and Transformer-based models are specific implementations of UAT, both exhibiting the characteristic of dynamically approximating functions based on inputs. Larger models and datasets are indeed requirements of the UAT theory: deeper network layers can more precisely approximate target functions dynamically based on inputs, while larger and more diverse datasets help ensure the training data closely represents real-world conditions, making the model’s approximated functions more reflective of real-world complexities.
However, this trend inevitably leads to increased demands for computational resources and extended training times. As more powerful models are needed, the UAT theory mathematically necessitates larger networks, which in turn require more computational resources. While technological advancements, especially in hardware performance, provide potential ways to alleviate this issue, the increase in time costs remains a fundamental obstacle to scaling serial networks. The improvement in computational resources such as GPUs and TPUs is limited by the speed of electrical signal transmission, whereas the theoretical scale of networks has no such upper limit. When networks reach tens or even hundreds of thousands of layers, inference times can extend to several minutes or even hours for each batch. Current solutions include model optimization techniques like quantization and pruning (Sun et al. 2023; Ma, Fang, and Wang 2023; Xia et al. 2023), hardware acceleration with specialized accelerators and edge computing (Gale et al. 2020), and parallel and distributed computing (Ben-Nun and Hoefler 2018) with model parallelization, data parallelization, and distributed computing frameworks. However, these do not fundamentally address the inference delay caused by the increasing number of serial network layers.
We believe the root cause lies in the serial computation mechanism of existing deep learning models, where the output of one layer serves as the input for the next. This design stems from the early development of deep learning networks in the field of CV, where image data exhibit local correlations and spatiotemporal invariance. To accommodate these features, early CV networks primarily employed convolutional layers with small receptive fields. To achieve a larger receptive field, deeper network layers were necessary, leading to early deep-learning architectures like AlexNet and VGG. The introduction of ResNet brought a paradigm shift, as residual networks have the ability to dynamically fit functions based on input, influencing almost all subsequent network designs and solidifying the tradition of multi-layer stacking.
Figure 1: The general description of the serial network and parallel network.
While this design was manageable with fewer layers, the increasing depth and number of parameters in modern networks have created a critical issue: slow computation efficiency due to the need for sequential layer-by-layer calculations during inference. This computation method inevitably leads to delays, which become more severe as the network depth increases, significantly constraining the progress of CV and NLP technologies.
Therefore, it is imperative to explore parallelization techniques for deep learning networks, enabling parallel computation across multiple layers. This approach is key to accelerating the inference process, breaking the time bottleneck, and advancing CV and NLP technologies towards greater efficiency and speed. To develop such parallel computation, we must start from the fundamental theory of deep learning: the UAT. The network must be designed in accordance with UAT principles and must satisfy the requirements outlined in the papers UAT2LLM and UAT2CV, which emphasize the ability to dynamically fit functions based on data. With this in mind and leveraging existing network structures, we designed a parallel network, Para-Former, to validate the feasibility of our theory. Our contributions include:
  • Proposing the theory of parallel deep learning networks and experimentally validating its feasibility.
  • Analyzing the characteristics, advantages, and disadvantages of both serial and parallel networks.
  • Investigating the issues present in deep learning from a data perspective.
The structure of this paper is as follows: In Section 2, we first present the general form of the Universal Approximation Theorem (UAT). Based on the mathematical form of UAT, we define the parallel network Para-Former in Section 3. In Section 4, we validate the effectiveness of Para-Former in various ways: we assess its feasibility in Section 4.2, explore the impact of network depth in Section 4.3, and analyze the influence of different datasets in Section 4.4.
2 UAT
Deep learning networks have demonstrated their powerful capabilities in numerous CV and NLP tasks. In the papers UAT2LLMs and UAT2CVs, it has been proven that both Transformer-based networks and residual-based CNNs are specific implementations of the UAT. Therefore, we consider UAT to be the fundamental theory for deep learning. To design parallelized networks, we must start from the principles of UAT. Before formally designing parallel networks, we will first review the basic form of the UAT theory and then propose a network parallelization solution based on this mathematical theory. The UAT was initially proposed by Cybenko (Cybenko 2007), and its basic form is as follows:
$G(\mathbf{x})=\sum_{j=1}^{N}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$ (1)
is dense in $C\left(\mathbf{I}_{n}\right)$. Here, $\mathbf{W}_{j}\in\mathbb{R}^{n}$ and $\alpha_{j},\theta\in\mathbb{R}$ are parameters $\sigma$ is a sigmoid function. For any $f\in C\left(\mathbf{I}_{n}\right)$ and $\varepsilon>0$, there exists a function $G(\mathbf{x})$:
$|G(\mathbf{x})-f(\mathbf{x})|<\varepsilon\quad\text{ for all }\quad\mathbf{x}\in\mathbf{I}_{n}.$ (2)
This indicates that when $N$ is sufficiently large, a neural network can approximate any continuous function on a closed interval. Hornik, Stinchcombe, and White  (Hornik, Stinchcombe, and White 1989) further proved that this function can approximate any Borel measurable function. Observing Equation (1), where the function $G(\mathbf{x})$ produces scalar outputs in the real number field $\mathbb{R}$, the scenario naturally extends when $G(\mathbf{x})$ maps to $\mathbb{R}^{m}$, requiring approximation in each dimension. To accommodate such multi-dimensional outputs, we simply need to extend Equation (1): the transformation matrix $\mathbf{W}_{j}$ is revised to the space $\mathbb{R}^{n\times m}$, the bias term $\theta_{j}$ is recast as a vector in $\mathbb{R}^{m}$, and $\alpha_{j}$ is reshaped into a matrix. We call the parameters $\mathbf{W}_{j}$ and $\alpha_{j}$ in UAT as weight and $\theta_{j}$ as bias.
Figure 2: The process of Transformer module.
3 UAT for Parallel Network
3.1 Pre-request for Basic Block of Parallel Network
From Equation (1), we can observe a significant characteristic: if we define different indices $j$ as the indices of layers and the corresponding layer’s mathematical form is $\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$, it is evident that the parameters between different layers do not interact with each other. Therefore, based on (1), we can design networks where parameters across multiple layers are independent, achieving parallel computation, like Figure 1. Figure 1.a represents a typical sequential network, which clearly requires computation to proceed in order from front to back. In contrast, Figure 1.b depicts a parallel network, where there is no parameter interaction between different blocks, allowing them to be considered independent and thus suitable for parallel computation. The blocks in this network should satisfy the formula given in Figure 1.b. Clearly, under this condition, the sum of the multiple blocks’ outputs adheres to the UAT theory.
If we design the parallel network solely based on Eq. (1), the network will face a limitation once training is complete: the function that can be approximated by the network is fixed, as the parameters in the corresponding UAT are fixed. So to ensure the network can dynamically approximate functions based on input, some or all of the parameters $\alpha_{j}$, $W_{j}$ and $\theta_{j}$ must be influenced by the input (referencing to UAT2LLM and UAT2CV), which can be simply written as $\alpha_{j}(\mathbf{x})$, $W_{j}(\mathbf{x})$ and $\theta_{j}(\mathbf{x})$.
This allows the network to dynamically adapt and fit the function based on the input while enabling parallel computation. So we could design the basic block in Figure 1, like $\text{Block}_{1}\cdots\text{Block}_{n}$ are the implementation of $\alpha_{j}(\mathbf{x}))\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}(\mathbf{x})\mathbf{x}+\theta_{j}(\mathbf{x})\right),i=1\cdots n$. Some or all the parameters $\alpha_{j}$, $W_{j}$ and $\theta_{j}$ will change according to input.
Figure 3: The structure of the Para-Former network. a.Para-Former-1-n and b.Para-Former-m-n represent Para-Former with 1 depth of n layers and m depth of n layers, respectively.
3.2 Para-former
Based on the previous request on the basic module, we designed the Para-Former which is a parallel network by using the Transformer module as the basic module. We chose Transformer for this design because it is a well-established and widely-used network architecture, excelling in both CV and NLP. Another reason is that its basic mathematical form aligns with the requirements of Section 3.1.
Before formally introducing the structure of Para-Former, let’s briefly review the Transformer structure. The basic structure of a Transformer is shown in Figure 2, and for convenience, we refer to it as a Transformer module. Typically, it is represented as follows:
$\mathring{\mathbf{H}}=$ $\operatorname{MultiHead}(\mathbf{Q},\mathbf{K},\mathbf{V})$ (3)
$=$ $\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}$
$\hat{\mathbf{H}}_{i}$ $=\operatorname{Attention}(\mathbf{x}_{i}\mathbf{W}_{Qi}=\mathbf{Q}_{i},\mathbf{x}_{i}\mathbf{W}_{Ki}=\mathbf{K}_{i},$ (4)
$\mathbf{x}_{i}\mathbf{W}_{Vi}=\mathbf{V}_{i})$
$=\operatorname{softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{T}}{\sqrt{M}}\right)\mathbf{V}_{i}$
$=\mathbf{H}_{i}\mathbf{V}_{i}=\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$
$FFN(\mathbf{x}_{i})=\mathbf{W}_{i,3}\sigma(\mathbf{W}_{i,2}\mathbf{x}_{i}+\mathbf{b}_{i,2})+\mathbf{b}_{i,3}\$ (5)
The above representation mainly describes the engineering process rather than its true mathematical form. According to UAT2LLM, the Multi-Head Attention (MHA) and Feed-Forward Network (FFN) components of Transformers can be expressed in matrix-vector form. We denote the matrix-vector form of variables with a prime symbol (e.g., $\mathbf{x}_{i}$ becomes $\mathbf{x}^{\prime}_{i}$). For more details, refer to UAT2LLM.
$MHA(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}$ (6)
$FFN(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}\$ (7)
The parameter $\mathbf{W}^{\prime}_{i,1}$ is influenced by the input. As shown in Figure 2, the output of a single Transformer layer, $\mathbf{x}^{\prime}_{i+1}$, can be described by Eq. (8) (Details could be found in Section A):
$\mathbf{x}^{\prime}_{i+1}$ $=\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}$ (8)
The mathematical form in Eq. (8) meets the requirements of a basic module in UAT. Therefore, we can use the structure in Figure 2 as the basic module in Para-Former, corresponding to Figure 3.a. Besides, we add a symbol of a hat on some parameters to represent the parameters will be affected by the input, like $\hat{\mathbf{W}^{\prime}_{i,1}}$.
$\mathbf{x}^{\prime}_{i+2}$ $=\hat{\mathbf{W}^{\prime}_{i+1,2}}\mathbf{x}^{\prime}_{i}+\hat{\mathbf{W}^{\prime}_{i+1,3}}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})$ (9)
$+\mathbf{W}^{\prime}_{i+1,3}\sigma\{\hat{\mathbf{W}^{\prime}_{i+1,4}}\mathbf{x}^{\prime}_{i}+\hat{\mathbf{b}^{\prime}_{i+1,4}}\}+\hat{\mathbf{b}^{\prime}_{i+1,5}}$
The output of a two-layer Transformer is shown in Eq. (9) (Details could be found in Section A). From Eq. (9), it is evident that a serial configuration of multi-layer Transformers also conforms to the mathematical form of UAT. Therefore, we can extend the Para-Former structure. That is, within one block in Figure 1, there can be multiple serially connected Transformer modules, while multiple blocks operate in parallel, as shown in Figure 3.b. We refer to the number of Transformer modules within a block as the depth and the total number of Transformer modules across all blocks as the network layer count.
3.3 Speed-up Ratio
We have proposed a network parallelization scheme, and in this section, we will compare the running speed of this parallelized network with that of traditional serial networks. It is important to note that due to differences in hardware and implementation methods, providing an objective comparison through experiments can be challenging. Therefore, we will focus on a theoretical analysis. Based on the network structure shown in Figure 2, the inference time of Para-Former is only affected by its network depth.
Assume we use a Para-Former with a depth of $M$ and a layer count of $L$, compared to a traditional Transformer network with a depth of $N$ layers, where $L>>N$. In this case, the running speed of the parallel network would be $N/M$ times faster than that of the serial network. This is because the speed of the parallel network is solely related to the depth of Para-Former. As the network depth increases, the acceleration effect becomes more pronounced.
4 Experiments
Our experimental approach is to first verify the feasibility of the parallel network theory and its adherence to the UAT approximation properties using several commonly used public image classification datasets (Section 4.1). We then explore the impact of network depth on the UAT theory (Section 4.3), followed by an investigation into the influence of the dataset (Section 4.4).
Table 1: This table presents the sizes of the training, validation, and test sets for the datasets CIFAR-10, CIFAR-100, STL-10, OxfordIIITPet, and Flower-102. It also provides the number of categories (Classes) and the average number of images per category for each dataset (Per class).
Dataset Cifar-10 Cifar-100 STL-10 OxfordIIITPet Flower-102
Classes 10 100 10 37 102
Train set 47500 47500 4750 3,515 969
Val set 2500 47500 250 185 51
Test set 10000 10000 800 3669 6149
Per class 4750 4750 475 95 9.5
4.1 Dataset
In pursuit of a comprehensive understanding of the efficacy and practicality of parallel models, we meticulously selected several widely-acknowledged and representative image classification datasets for an integrated comparative analysis. These datasets encompass a diverse array of visual content and span various levels of difficulty, thereby furnishing us with a fertile ground to probe the performance boundaries of parallel models. Specifically, our study centers on the following pivotal datasets:
1. CIFAR-10 & CIFAR-100 (Krizhevsky 2009): These were constructed by Alex Krizhevsky et al. CIFAR-10 encompasses 60,000 32x32 pixel color images across 10 classes, while CIFAR-100 extends this to 100 categories.
2. STL-10 (Coates, Ng, and Lee 2011): STL-10 comprises 5,000 training and 8,000 testing images distributed among 10 classes, each at a 96x96 pixel resolution.
3. Flowers-102 (Nilsback and Zisserman 2008): Focused on floral categorization, this dataset houses 102 distinct flower species, totalling 8,189 images with a per-species count ranging from 40 to 258.
4. Oxford-IIIT Pets (Parkhi et al. 2012): Tailored for the recognition of pet breeds (cats and dogs), this dataset contains over 7,300 images across 37 categories, characterized by varied animal poses, backgrounds, and lighting conditions.
In Table 1, we provide more details for these dataset.
Through experimentation on these meticulously chosen datasets, we aspire to thoroughly validate the feasibility and strengths of parallel models across a spectrum of scales, complexities, and domains in image classification tasks. Simultaneously, we seek to elucidate potential limitations and chart avenues for future research endeavors.
Table 2: This table shows the predictions of Para-former-1-1, Para-former-1-6, Para-former-1-12, Para-former-1-24 on the datasets CIFAR-10, CIFAR-100, STL-10, OxfordIIITPet, and Flower-102. We simply represent them as Para-F-1-1, Para-F-1-6, Para-F-1-12 and Para-F-1-24, simply.
Dataset Para-F-1-1 Para-F-1-6 Para-F-1-12 Para-F-1-24
Cifar-10 55.95 64.90 66.88 71.53
Cifar-100 30.67 38.84 41.41 46.96
STL-10 44.41 51.57 55.96 59.53
OxfordIIITPet 19.62 20.95 22.07 22.73
Flower-102 27.33 32.65 37.42 35.84
4.2 The Feasibility of Para-Former
To validate the feasibility and characteristics of Para-Former, we conducted training and validation on several commonly used image classification datasets, with all models trained for 500 epochs. The experimental design involved preprocessing inputs using the common ViT techniques of patch extraction and positional encoding. We then used Transformer as the basic model architecture, setting all network depths to 1 and configuring Para-Former from single-layer to multi-layer setups, as shown in Figure 3.a. A single-layer setup corresponds to a ViT with a depth of 1. The results of the models are presented in Table 2.
It is evident from Table 2 that as the number of parallel network layers increases, the model accuracy improves progressively (Except for the result of Para-F-1-24 being worse than Para-F-1-12 on the Flower-102 dataset, this variation is within a reasonable range.). According to UAT theory, increasing the network layers is equivalent to increasing $N$ in Equation (1), thereby enhancing the approximation capability of UAT. The results in Table 2 are in complete agreement with UAT theory. Therefore, we believe that this parallel network approach is entirely feasible.
However, the overall prediction accuracy in Table 2 is quite low. What could be the reasons for this? We believe there are three main factors affecting the model’s predictive performance:
1. Model’s Fitting Capability: This is influenced by the model’s size and design. Specifically, the value of $N$ in UAT and the degrees of freedom in the parameters. The larger the $N$ value, the stronger the UAT’s fitting capability. Parameter freedom refers to the number of parameters affected by the input in UAT, such as a freedom of 1 in Eq. 8 and a freedom of 6 in Eq. 9. The degrees of freedom correspond to the model’s ability to dynamically fit based on input. Merely increasing $N$ enhances the model’s ability to fit specific functions, but images are often diverse. Therefore, the network must have the ability to dynamically fit the corresponding function based on the input. Higher degrees of freedom mean the model can fit a wide range of features with significant differences. Additionally, the bias term in the network can also dynamically fit based on input using UAT, which we refer to as bias freedom. Models with bias freedom have a stronger dynamic fitting capability. The mathematical model generated by the residual network is an example of UAT with bias freedom, achieving the fitting capability of a multi-layer network with just shallow layers. We will talk about this problem in Section 4.3.
2. Data Influence: Data is another critical factor in deep learning training. Since natural images are usually diverse, limited data can lead to overfitting. For example, in a cat-dog classification task, if dogs in the training set mostly appear against backgrounds of blue skies and grass, these backgrounds might be mistakenly considered part of the dog’s features, a phenomenon known as feature coupling. If cats rarely appear in such backgrounds in the training set but do in the test set, they might be misclassified as dogs. Training on a large, diverse dataset means that common backgrounds frequently appear with various objects or animals, such as cats, dogs, and elephants against blue skies, white clouds, houses, and trees. Thus, even when classifying objects, the model inherently learns background information. In this context, fine-tuning on the cat-dog dataset yields better results because the background does not interfere with dog feature recognition. Without sufficient data, even a network with strong fitting capabilities cannot perform well and is prone to overfitting. This problem will be considered in Section 4.4.
Considering these factors, it is clear that simply improving the model architecture is not enough. A robust training dataset is equally important for achieving high predictive accuracy.
Table 3: This table shows the predictions of Para-former-6-6, Para-former-2-24, Para-former-3-24, Para-former-6-24 on the datasets CIFAR-10, CIFAR-100, STL-10, OxfordIIITPet, and Flower-102.
Dataset Para-F-6-6 Para-F-2-24 Para-F-3-24 Para-F-6-24
Cifar-10 79.39 74.75 77.21 79.92
Cifar-100 55.18 51.04 52.65 56.15
STL-10 62.81 61.60 62.75 64.30
OxfordIIITPet 26.05 23.65 26.02 25.92
Flower-102 37.97  39.17 40.12 40.51
4.3 Model Depth Effect
Based on the conclusions from Section 4.2, we will explore the impact of network depth on predictive performance in this section. According to Eq. 8 and Eq. 9, we know that model depth primarily affects the degrees of freedom of the model parameters in UAT. Therefore, we designed Para-Former models with 24 layers and varying depths of 2, 3, and 6, and compared their predictive performance with that of a ViT with a depth of 6. The results are shown in Table 3.
It is evident that all results surpass those of Para-F-1-24 in Table 2 because the networks in Table 3 have higher degrees of freedom in their parameters. Notably, Para-F-6-6 consistently outperforms Para-F-2-24 and Para-F-3-24, indicating that the higher the degrees of freedom of the parameters, the stronger the model’s performance. However, the model cannot be too deep, so we must balance network depth and layer count. Another issue is that almost all models hit a performance bottleneck. Particularly, the prediction accuracy for OxfordIIITPet and Flower-102 is very poor. This is because these datasets are relatively small and exhibit high data diversity.
Figure 4: OxfordIIITPet dataset’s accuracy on different classes.
Figure 5: Flower-102 dataset’s accuracy on different classes.
4.4 Dataset Effect
Table 4: This table shows the results of fine-tuning using Google’s official pre-trained parameters. The pre-trained parameters were obtained from training the ViT model on the ImageNet dataset.
Dataset Cifar-10 Cifar-100 STL-10 OxfordIIITPet Flower-102
Para-F-12-12 98.62 91.36 99.47 91.87 96.89
Current deep learning models generally perform well on the aforementioned datasets. However, as shown in Table 3, the prediction accuracy of Para-Former is not very high, especially on the OxfordIIITPet and Flower-102 datasets. What is the fundamental reason for this poor performance? In this section, we will explain the relationship between deep learning and data.
First, the results were obtained using the validated ViT model (Para-F-1-1 and Para-Former-6-6), which performed similarly to Para-Former. Given that the ViT model has been extensively validated and is a commonly used model architecture, the effectiveness of the models is unquestionable. So, why is the performance so poor? The root cause lies in the data itself: the model’s performance is generally directly proportional to the amount of available data. Take the Flower-102 dataset, for example, which includes 102 classes of flowers but has only 1,020 images in the training set, with some reserved for validation. This is far from sufficient to train deep learning models. Why, then, does the OxfordIIITPet dataset, which has more data per class on average than Flower-102, perform even worse? This is due to the diversity of the data (i.e., the similarity of features within the same class) and the variety of backgrounds.
Figures 4 and 5 show some results from the OxfordIIITPet and Flower-102 datasets. Both figures indicate that prediction results are worse for categories with diverse backgrounds or inherent diversity, as seen in Figures 4.b and 5.b. Conversely, results are relatively better when the background is consistent but the species vary, as shown in Figures 4.a and 5.a. The best results are achieved when both the background and species variation are minimal, as shown in Figures 4.c and 5.c. Although the Flower-102 dataset has fewer images, its backgrounds are more consistent and there is less variation among the species. In contrast, animals in the OxfordIIITPet dataset can appear in various forms, while flowers in the Flower-102 dataset are usually photographed from relatively fixed angles, leading to better performance.
Does this mean that deep learning models are not powerful enough or lack generalization and reasoning capabilities? On the contrary, achieving around 25% accuracy on OxfordIIITPet and about 40% on Flower-102 with such limited data actually demonstrates the strength and generalization capability of deep learning models. A common misconception about deep learning is that humans, even if unfamiliar with certain species, can accurately distinguish OxfordIIITPet and Flower-102 images by carefully examining them, thus believing that deep learning lacks generalization or human-like inferencing ability. This viewpoint is flawed. Assuming the human brain functions as a UAT model, it can automatically disregard background information before seeing these images, as this background is frequently encountered in daily life, and the UAT in the human brain has already fitted this information. Therefore, humans only need to focus on specific information and then use it to train their brains to find similarities and differences. For instance, a 10-year-old child’s brain, observing the world at 60 frames per second, would have processed about $1\times 10^{10}$ high-resolution images by the age of 10, many of which include common objects and backgrounds in daily life. Therefore, criticizing deep learning models for being unable to train on small datasets and lacking universality and generalization is unfair.
Thus, we should train on large datasets and then fine-tune on smaller ones. This is also the current deep learning training method, showcasing the powerful generalization capability of deep learning models, especially when there is some correlation between datasets. Following this fine-tuning approach, we fine-tuned the pre-trained ViT model on the aforementioned datasets for just 10 epochs. The results, as shown in Table 4, indicate a significant improvement in prediction performance.
Figure 6 shows the rationality behind fine-tuning. Since we essentially adjust parameters based on the input to produce target results using UAT, if there is a correlation in the data, the corresponding parameters derived from the input ($\mathbf{\alpha}_{i},\mathbf{W}_{i}$, and $\mathbf{\theta}_{i};i=1,\cdot,N$) will be similar. Therefore, fine-tuning is feasible.
Figure 6: The fine-tuning process from the perspective of UAT.
5 Conclusion
In this paper, we propose a parallel network based on the Universal Approximation Theorem (UAT) and validate its effectiveness on large datasets. We analyze the characteristics of both serial and parallel networks. Serial networks, especially those based on residual connections, often have higher degrees of parameter freedom, leading to stronger generalization capabilities. However, as the network depth increases, the inference time also becomes longer. On the other hand, parallel networks trade computational resources for time, meaning their runtime is not affected by the number of network layers. This allows for the design of multi-layer parallel networks that can replace serial networks to achieve fast inference. Additionally, we analyze the impact and reasons data have on deep learning. From the UAT perspective, training on large datasets essentially enhances the model’s ability to fit data.
Appendix A The UAT Format of Transformer
A.1 The UAT Format of Transformer-based ViTs
Eq.(1) and (2) give an example of the matrix-vector format of MHA and FFN. Eq. (3) and Eq. (4) give the corresponding UAT format of $\mathbf{x}_{i+1}$ and $\mathbf{x}_{i+2}$. It is obvious that the bias and weight terms in the UAT corresponding to Transformed can change dynamically based on the input.
Due to the complexity of the variables involved, we will combine some terms to align the resulting equations with the standard UAT form. To distinguish these parameter variables, we will keep the original variables unchanged. If a parameter variable is influenced only by other variables, we will denote it with a bar on top, such as $\bar{\mathbf{W}}$. If it is influenced by the input $\mathbf{x}_{i}$, we will mark it with a hat, such as $\hat{\mathbf{W}}$.
$MHA(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}$ (1)
$FFN(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}\$ (2)
$\mathbf{x}^{\prime}_{i+1}$ $=\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma[\mathbf{W}^{\prime}_{i,2}(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i})+\mathbf{b}^{\prime}_{i,2}]+\mathbf{b}^{\prime}_{i,3}$ (3)
$=\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma(\underline{\mathbf{W}^{\prime}_{i,2}\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}$
$=\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}$
$\mathbf{x}^{\prime}_{i+2}$ $=\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i+1}+\mathbf{W}^{\prime}_{i+1,3}\sigma(\hat{\mathbf{W}^{\prime}_{i+1,1}}\mathbf{x}^{\prime}_{i+1}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{b}^{\prime}_{i+1,3}$ (4)
$=\mathbf{W}^{\prime}_{i+1,1}[\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}]$
$+\mathbf{W}^{\prime}_{i+1,3}\sigma\{\hat{\mathbf{W}^{\prime}_{i+1,1}}[\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}]+\mathbf{b}^{\prime}_{i+1,2}\}+\mathbf{b}^{\prime}_{i+1,3}$
$=\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}^{\prime}_{i,3}}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{b}^{\prime}_{i,3}}$
$+\mathbf{W}^{\prime}_{i+1,3}\sigma\{\underline{\hat{\mathbf{W}^{\prime}_{i+1,1}}\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\underline{\hat{\mathbf{W}^{\prime}_{i+1,1}}\mathbf{W}^{\prime}_{i,3}}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\underline{\hat{\mathbf{W}^{\prime}_{i+1,1}}\mathbf{b}^{\prime}_{i,3}+\mathbf{b}^{\prime}_{i+1,2}}\}+\mathbf{b}^{\prime}_{i+1,3}$
$=\hat{\mathbf{W}^{\prime}_{i+1,2}}\mathbf{x}^{\prime}_{i}+\hat{\mathbf{W}^{\prime}_{i+1,3}}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{W}^{\prime}_{i+1,3}\sigma\{\hat{\mathbf{W}^{\prime}_{i+1,4}}\mathbf{x}^{\prime}_{i}+\underline{\hat{\mathbf{W}^{\prime}_{i+1,5}}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\hat{\mathbf{b}^{\prime}_{i+1,2}}}\}+\underline{\hat{\mathbf{b}^{\prime}_{i,3}}+\mathbf{b}^{\prime}_{i+1,3}}$
$=\hat{\mathbf{W}^{\prime}_{i+1,2}}\mathbf{x}^{\prime}_{i}+\hat{\mathbf{W}^{\prime}_{i+1,3}}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{W}^{\prime}_{i+1,3}\sigma\{\hat{\mathbf{W}^{\prime}_{i+1,4}}\mathbf{x}^{\prime}_{i}+\hat{\mathbf{b}^{\prime}_{i+1,4}}\}+\hat{\mathbf{b}^{\prime}_{i+1,5}}$

### v4

@@
 Neural networks are increasingly evolving towards training large models with big data, a method that has demonstrated superior performance across many tasks.
-However, this approach introduces an urgent problem: current deep learning models are predominantly serial, meaning that as the number of network layers increases, so do the training and inference times.
-This is unacceptable if deep learning is to continue advancing.
+However, this approach introduces an urgent problem: current deep learning models are serial computing, meaning that as the number of network layers increases, so do the training and inference times.
+This is unacceptable if deep learning models go deeper.
 Therefore, this paper proposes a deep learning parallelization strategy based on the Universal Approximation Theorem (UAT).
@@
 1 Introduction
-In today’s era, deep learning models have undoubtedly taken the lead in the fields of computer vision (CV) and natural language processing (NLP), demonstrating exceptional performance across numerous tasks.
-In CV, models such as ResNet(He et al.
-2015) and ViT(Vaswani et al.
-2017) for image recognition, U-Net (Ronneberger, Fischer, and Brox 2015) for image segmentation, GAN (Goodfellow et al.
-2014) for image generation, and diffusion models like SORA (Liu et al.
-2024) for video generation highlight the extensive application of deep learning in CV.
-These models are primarily based on convolutional neural networks (CNNs), Transformers, or a combination of both.
-Simultaneously, large language models (LLMs) based on Transformers have become dominant in natural language processing.
-Overall, deep learning models have shown vast application potential and powerful capabilities so far.
-However, this progress has brought forth some challenges: the development trend in deep learning is moving towards larger models and larger datasets.
-This trend is fundamentally driven by the UAT(Cybenko 2007; Hornik, Stinchcombe, and White 1989).
-From the papers UAT2LLM(Wang and Li 2024) and UAT2CV(Wang and Li 2024), we know that residual-based CNNs and Transformer-based models are specific implementations of UAT, both exhibiting the characteristic of dynamically approximating functions based on inputs.
-Larger models and datasets are indeed requirements of the UAT theory: deeper network layers can more precisely approximate target functions dynamically based on inputs, while larger and more diverse datasets help ensure the training data closely represents real-world conditions, making the model’s approximated functions more reflective of real-world complexities.
-However, this trend inevitably leads to increased demands for computational resources and extended training times.
-As more powerful models are needed, the UAT theory mathematically necessitates larger networks, which in turn require more computational resources.
-While technological advancements, especially in hardware performance, provide potential ways to alleviate this issue, the increase in time costs remains a fundamental obstacle to scaling serial networks.
-The improvement in computational resources such as GPUs and TPUs is limited by the speed of electrical signal transmission, whereas the theoretical scale of networks has no such upper limit.
-When networks reach tens or even hundreds of thousands of layers, inference times can extend to several minutes or even hours for each batch.
-Current solutions include model optimization techniques like quantization and pruning (Sun et al.
-2023; Ma, Fang, and Wang 2023; Xia et al.
-2023), hardware acceleration with specialized accelerators and edge computing (Gale et al.
-2020), and parallel and distributed computing (Ben-Nun and Hoefler 2018) with model parallelization, data parallelization, and distributed computing frameworks.
-However, these do not fundamentally address the inference delay caused by the increasing number of serial network layers.
-We believe the root cause lies in the serial computation mechanism of existing deep learning models, where the output of one layer serves as the input for the next.
-This design stems from the early development of deep learning networks in the field of CV, where image data exhibit local correlations and spatiotemporal invariance.
-To accommodate these features, early CV networks primarily employed convolutional layers with small receptive fields.
-To achieve a larger receptive field, deeper network layers were necessary, leading to early deep-learning architectures like AlexNet and VGG.
-The introduction of ResNet brought a paradigm shift, as residual networks have the ability to dynamically fit functions based on input, influencing almost all subsequent network designs and solidifying the tradition of multi-layer stacking.
-Figure 1: The general description of the serial network and parallel network.
+Currently, deep learning is advancing towards larger models and datasets, driven by the success of ChatGPT (Radford and Narasimhan 2018; Brown et al.
+2020; Achiam et al.
+2023), Llama (Touvron et al.
+2023), and ViT (Dosovitskiy et al.
+2020), among others.
+These large models demonstrate significantly enhanced performance compared to their predecessors.
+According to UAT2LLMs (Wang and Li 2024b) and UAT2CVs (Wang and Li 2024b), residual-based convolutional neural networks (CNNs) and Transformer-based large language models are specific implementations of the UAT (Cybenko 2007; Hornik, Stinchcombe, and White 1989), capable of dynamically approximating functions based on inputs.
+From the perspective of UAT, larger models inherently possess stronger dynamic fitting capabilities (Wang and Li 2024b, a).
+Therefore, the trend towards larger models is undoubtedly the future direction of deep learning, and large datasets are essential for training these models to ensure their approximations accurately reflect real-world complexities.
+However, this trend towards larger models and datasets inevitably leads to increased demands for computational resources and extended training times.
+Despite advancements in hardware technology, such as GPUs and TPUs, the rise in time costs remains a significant obstacle.
+The speed of electrical signal transmission is limited, whereas the theoretical scale of networks has no such upper limit.
+As network layers increase to tens or even hundreds of thousands, inference times can extend significantly, potentially requiring several minutes or hours for a single batch, making training large models on large datasets exceedingly difficult.
+Current solutions include model optimization techniques such as quantization (Jacob et al.
+2017; Lu, Zhong, and Yang 2022; Guo, Wang, and Cui 2021; Lin et al.
+2023), pruning (Ma, Fang, and Wang 2023; Xia et al.
+2023), and knowledge distillation (Gou et al.
+2020; Gu et al.
+2021).
+Additionally, specialized accelerators (Jouppi et al.
+2017; Elbtity, Chandarana, and Zand 2024) and edge computing (Shi et al.
+2016; Gale et al.
+2020), as well as parallel and distributed computing (Shallue et al.
+2018; Ben-Nun and Hoefler 2018; Sergeev and Balso 2018), offer potential solutions.
+However, these methods do not fundamentally address the inference delay caused by the increasing number of serial network layers.
+The root cause lies in the serial computation mechanism of existing deep learning models, where the output of one layer serves as the input for the next.
+This design originated from early computer vision networks, which used convolutional layers with small receptive fields to accommodate the local correlations and spatiotemporal invariance of image data.
+Achieving a larger receptive field required deeper network layers, as exemplified by architectures like AlexNet (Krizhevsky, Sutskever, and Hinton 2012) and VGG (Simonyan and Zisserman 2014).
+The introduction of ResNe t (He et al.
+2015) marked a paradigm shift, as the residual structure could dynamically fit functions based on input (Wang and Li 2024b, a), influencing subsequent network designs and establishing the tradition of multi-layer stacking.
 While this design was manageable with fewer layers, the increasing depth and number of parameters in modern networks have created a critical issue: slow computation efficiency due to the need for sequential layer-by-layer calculations during inference.
-This computation method inevitably leads to delays, which become more severe as the network depth increases, significantly constraining the progress of CV and NLP technologies.
+This computation method inevitably leads to delays, which become more severe as network depth increases, significantly constraining the progress of computer vision and natural language processing technologies.
 Therefore, it is imperative to explore parallelization techniques for deep learning networks, enabling parallel computation across multiple layers.
-This approach is key to accelerating the inference process, breaking the time bottleneck, and advancing CV and NLP technologies towards greater efficiency and speed.
+This approach is key to accelerating the inference process, breaking the time bottleneck, and advancing computer vision and natural language processing technologies towards greater efficiency and speed.
 To develop such parallel computation, we must start from the fundamental theory of deep learning: the UAT.
-The network must be designed in accordance with UAT principles and must satisfy the requirements outlined in the papers UAT2LLM and UAT2CV, which emphasize the ability to dynamically fit functions based on data.
+The network must be designed in accordance with UAT principles and must satisfy the requirements outlined in the papers UAT2LLMs (Wang and Li 2024b) and UAT2CVs (Wang and Li 2024a), which emphasize the ability to dynamically fit functions based on data.
 With this in mind and leveraging existing network structures, we designed a parallel network, Para-Former, to validate the feasibility of our theory.
@@
   • Investigating the issues present in deep learning from a data perspective.
-The structure of this paper is as follows: In Section 2, we first present the general form of the Universal Approximation Theorem (UAT).
+The structure of this paper is as follows: In Section 2, we first present the general form of the UAT.
 Based on the mathematical form of UAT, we define the parallel network Para-Former in Section 3.
-In Section 4, we validate the effectiveness of Para-Former in various ways: we assess its feasibility in Section 4.2, explore the impact of network depth in Section 4.3, and analyze the influence of different datasets in Section 4.4.
+In Section 4, we validate the effectiveness of Para-Former in various ways: we assess its feasibility in Section 4.2, explore the impact of network depth in Section 4.3, and analyze the influence of datasets on network performances in Section 4.4.
 2 UAT
 Deep learning networks have demonstrated their powerful capabilities in numerous CV and NLP tasks.
-In the papers UAT2LLMs and UAT2CVs, it has been proven that both Transformer-based networks and residual-based CNNs are specific implementations of the UAT.
+In the papers UAT2LLMs (Wang and Li 2024b) and UAT2CVs (Wang and Li 2024a), they have been proven that both Transformer-based networks and residual-based CNNs are specific implementations of the UAT.
 Therefore, we consider UAT to be the fundamental theory for deep learning.
@@
 Hornik, Stinchcombe, and White  (Hornik, Stinchcombe, and White 1989) further proved that this function can approximate any Borel measurable function.
-Observing Equation (1), where the function $G(\mathbf{x})$ produces scalar outputs in the real number field $\mathbb{R}$, the scenario naturally extends when $G(\mathbf{x})$ maps to $\mathbb{R}^{m}$, requiring approximation in each dimension.
-To accommodate such multi-dimensional outputs, we simply need to extend Equation (1): the transformation matrix $\mathbf{W}_{j}$ is revised to the space $\mathbb{R}^{n\times m}$, the bias term $\theta_{j}$ is recast as a vector in $\mathbb{R}^{m}$, and $\alpha_{j}$ is reshaped into a matrix.
+Observing Eq.
+(1), where the function $G(\mathbf{x})$ produces scalar outputs in the real number field $\mathbb{R}$, the scenario naturally extends when $G(\mathbf{x})$ maps to $\mathbb{R}^{m}$, requiring approximation in each dimension.
+To accommodate such multi-dimensional outputs, we simply need to extend Eq.
+(1): the transformation matrix $\mathbf{W}_{j}$ is revised to the space $\mathbb{R}^{n\times m}$, the bias term $\theta_{j}$ is recast as a vector in $\mathbb{R}^{m}$, and $\alpha_{j}$ is reshaped into a matrix.
 We call the parameters $\mathbf{W}_{j}$ and $\alpha_{j}$ in UAT as weight and $\theta_{j}$ as bias.
-Figure 2: The process of Transformer module.
+Figure 1: The general description of the serial network and parallel network.
 3 UAT for Parallel Network
-3.1 Pre-request for Basic Block of Parallel Network
-From Equation (1), we can observe a significant characteristic: if we define different indices $j$ as the indices of layers and the corresponding layer’s mathematical form is $\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$, it is evident that the parameters between different layers do not interact with each other.
-Therefore, based on (1), we can design networks where parameters across multiple layers are independent, achieving parallel computation, like Figure 1.
+3.1 Requirements for Basic Block in Parallel Network
+From Eq.
+(1), we can observe a significant characteristic: if we define different indices $j$ as the indices of layers and the corresponding layer’s mathematical form is $\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$, it is evident that the parameters between different layers do not interact with each other.
+Therefore, based on Eq.
+(1), we can design networks where parameters across multiple layers are independent, achieving parallel computation, like Figure 1.
 Figure 1.a represents a typical sequential network, which clearly requires computation to proceed in order from front to back.
 In contrast, Figure 1.b depicts a parallel network, where there is no parameter interaction between different blocks, allowing them to be considered independent and thus suitable for parallel computation.
-The blocks in this network should satisfy the formula given in Figure 1.b.
+The blocks in this network should satisfy the formula of $\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$.
 Clearly, under this condition, the sum of the multiple blocks’ outputs adheres to the UAT theory.
@@
 (1), the network will face a limitation once training is complete: the function that can be approximated by the network is fixed, as the parameters in the corresponding UAT are fixed.
-So to ensure the network can dynamically approximate functions based on input, some or all of the parameters $\alpha_{j}$, $W_{j}$ and $\theta_{j}$ must be influenced by the input (referencing to UAT2LLM and UAT2CV), which can be simply written as $\alpha_{j}(\mathbf{x})$, $W_{j}(\mathbf{x})$ and $\theta_{j}(\mathbf{x})$.
+So to ensure the network can dynamically approximate functions based on input, some or all of the parameters $\alpha_{j}$, $W_{j}$ and $\theta_{j}$ must be influenced by the input (referencing to UAT2LLMs and UAT2CVs), which can be simply written as $\alpha_{j}(\mathbf{x})$, $W_{j}(\mathbf{x})$ and $\theta_{j}(\mathbf{x})$.
 This allows the network to dynamically adapt and fit the function based on the input while enabling parallel computation.
-So we could design the basic block in Figure 1, like $\text{Block}_{1}\cdots\text{Block}_{n}$ are the implementation of $\alpha_{j}(\mathbf{x}))\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}(\mathbf{x})\mathbf{x}+\theta_{j}(\mathbf{x})\right),i=1\cdots n$.
+So we could design the basic block in Figure 1, like $\text{Block}_{1}\cdots\text{Block}_{n}$ are the implementation of $\alpha_{j}(\mathbf{x})\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}(\mathbf{x})\mathbf{x}+\theta_{j}(\mathbf{x})\right),i=1\cdots n$.
 Some or all the parameters $\alpha_{j}$, $W_{j}$ and $\theta_{j}$ will change according to input.
-Figure 3: The structure of the Para-Former network. a.Para-Former-1-n and b.Para-Former-m-n represent Para-Former with 1 depth of n layers and m depth of n layers, respectively.
+Figure 2: The process of Transformer module.
+Figure 3: The structure of the Para-Former-1-n represents Para-Former with 1 depth of n layers.
+Figure 4: The structure of the Para-Former-m-n represents Para-Former with m depth of n layers.
+Cifar-10 Cifar-100 STL-10 OxfordIIITPet Flower-102
+Classes 10 100 10 37 102
+Train set 47500 47500 4750 3,515 969
+Val set 2500 47500 250 185 51
+Test set 10000 10000 800 3669 6149
+Per class 4750 4750 475 95 9.5
+Table 1: This table presents the sizes of the training, validation, and test sets for the datasets CIFAR-10, CIFAR-100, STL-10, OxfordIIITPet, and Flower-102.
+It also provides the number of categories (Classes) and the average number of images per category for each dataset (Per class).
 3.2 Para-former
@@
 $\mathring{\mathbf{H}}=$ $\operatorname{MultiHead}(\mathbf{Q},\mathbf{K},\mathbf{V})$ (3)
-$=$ $\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}$
-$\hat{\mathbf{H}}_{i}$ $=\operatorname{Attention}(\mathbf{x}_{i}\mathbf{W}_{Qi}=\mathbf{Q}_{i},\mathbf{x}_{i}\mathbf{W}_{Ki}=\mathbf{K}_{i},$ (4)
-$\mathbf{x}_{i}\mathbf{W}_{Vi}=\mathbf{V}_{i})$
-$=\operatorname{softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{T}}{\sqrt{M}}\right)\mathbf{V}_{i}$
-$=\mathbf{H}_{i}\mathbf{V}_{i}=\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$
+$=$ $\operatorname{Concat}\left(\hat{\mathbf{H}}_{i1},\ldots,\hat{\mathbf{H}}_{ih}\right)\mathbf{W}_{O}$
+$\hat{\mathbf{H}}_{ij}$ $=\operatorname{Attention}(\mathbf{x}_{ij}\mathbf{W}_{Qij}=\mathbf{Q}_{ij},\mathbf{x}_{ij}\mathbf{W}_{Kij}=\mathbf{K}_{ij},$ (4)
+$\mathbf{x}_{i}\mathbf{W}_{Vi}=\mathbf{V}_{ij})$
+$=\operatorname{softmax}\left(\frac{\mathbf{Q}_{ij}\mathbf{K}_{ij}^{T}}{\sqrt{M}}\right)\mathbf{V}_{ij}$
+$=\mathbf{H}_{ij}\mathbf{V}_{ij}=\mathbf{H}_{ij}[\mathbf{x}_{ij}\mathbf{W}_{Vij}]$
+where the subscript of $ij$ represents the $j$-th head of Multi-Head Attention (MHA) in the $i$-th layer.
+Please note that throughout the following text, all the subscript $i$ refer to the layer index in the network, like $\mathbf{x}_{i}$.
+According to Figure 2, assuming the input to the Feed-Forward Network (FFN) in the $i$-th layer of the Transformer, it can be expressed as:
 $FFN(\mathbf{x}_{i})=\mathbf{W}_{i,3}\sigma(\mathbf{W}_{i,2}\mathbf{x}_{i}+\mathbf{b}_{i,2})+\mathbf{b}_{i,3}\$ (5)
 The above representation mainly describes the engineering process rather than its true mathematical form.
-According to UAT2LLM, the Multi-Head Attention (MHA) and Feed-Forward Network (FFN) components of Transformers can be expressed in matrix-vector form.
-We denote the matrix-vector form of variables with a prime symbol (e.g., $\mathbf{x}_{i}$ becomes $\mathbf{x}^{\prime}_{i}$).
-For more details, refer to UAT2LLM.
-$MHA(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}$ (6)
-$FFN(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}\$ (7)
-The parameter $\mathbf{W}^{\prime}_{i,1}$ is influenced by the input.
-As shown in Figure 2, the output of a single Transformer layer, $\mathbf{x}^{\prime}_{i+1}$, can be described by Eq.
-(8) (Details could be found in Section A):
-$\mathbf{x}^{\prime}_{i+1}$ $=\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}$ (8)
-The mathematical form in Eq.
-(8) meets the requirements of a basic module in UAT.
-Therefore, we can use the structure in Figure 2 as the basic module in Para-Former, corresponding to Figure 3.a.
-Besides, we add a symbol of a hat on some parameters to represent the parameters will be affected by the input, like $\hat{\mathbf{W}^{\prime}_{i,1}}$.
-$\mathbf{x}^{\prime}_{i+2}$ $=\hat{\mathbf{W}^{\prime}_{i+1,2}}\mathbf{x}^{\prime}_{i}+\hat{\mathbf{W}^{\prime}_{i+1,3}}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})$ (9)
-$+\mathbf{W}^{\prime}_{i+1,3}\sigma\{\hat{\mathbf{W}^{\prime}_{i+1,4}}\mathbf{x}^{\prime}_{i}+\hat{\mathbf{b}^{\prime}_{i+1,4}}\}+\hat{\mathbf{b}^{\prime}_{i+1,5}}$
-The output of a two-layer Transformer is shown in Eq.
-(9) (Details could be found in Section A).
+According to UAT2LLMs, the MHA and FFN components of Transformers can be expressed in matrix-vector form.
+We denote the matrix-vector form of variables and operations with a prime symbol (e.g., $\mathbf{x}_{i}$ becomes $\mathbf{x}^{\prime}_{i}$).
+So MHA and FFN can be written as (For more details, refer to UAT2LLMs):
+$MHA^{\prime}(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}$ (6)
+$FFN^{\prime}(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{b}^{\prime}_{i+1,3}\$ (7)
+According to Eq.
+(6) and Eq.
+(7), the output $\mathbf{x}^{\prime}_{1}$ of a single-layer Transformer can be expressed as Eq.
+(8) (see Appendix A for more details).
+The mathematical form of Eq.
+(8) can be considered equivalent to the two terms $\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$ in Eq.
+(1).
+Therefore, if we parallelize multiple instances of $\mathbf{x}^{\prime}_{1}$, we obtain a multi-layer parallel Transformer network, as illustrated in Figure.
+3.
+$\mathbf{x}^{\prime}_{1}$ $=(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,2})$ (8)
+Additionally, we provide the output for a two-layer Transformer, as shown in Eq.
+(9) (see Appendix A for more details).
 From Eq.
-(9), it is evident that a serial configuration of multi-layer Transformers also conforms to the mathematical form of UAT.
+(9), it is evident that the serial configuration of two Transformer layers also adheres to the mathematical form of UAT.
 Therefore, we can extend the Para-Former structure.
-That is, within one block in Figure 1, there can be multiple serially connected Transformer modules, while multiple blocks operate in parallel, as shown in Figure 3.b.
-We refer to the number of Transformer modules within a block as the depth and the total number of Transformer modules across all blocks as the network layer count.
-3.3 Speed-up Ratio
-We have proposed a network parallelization scheme, and in this section, we will compare the running speed of this parallelized network with that of traditional serial networks.
-It is important to note that due to differences in hardware and implementation methods, providing an objective comparison through experiments can be challenging.
-Therefore, we will focus on a theoretical analysis.
-Based on the network structure shown in Figure 2, the inference time of Para-Former is only affected by its network depth.
-Assume we use a Para-Former with a depth of $M$ and a layer count of $L$, compared to a traditional Transformer network with a depth of $N$ layers, where $L>>N$.
-In this case, the running speed of the parallel network would be $N/M$ times faster than that of the serial network.
-This is because the speed of the parallel network is solely related to the depth of Para-Former.
-As the network depth increases, the acceleration effect becomes more pronounced.
-4 Experiments
-Our experimental approach is to first verify the feasibility of the parallel network theory and its adherence to the UAT approximation properties using several commonly used public image classification datasets (Section 4.1).
-We then explore the impact of network depth on the UAT theory (Section 4.3), followed by an investigation into the influence of the dataset (Section 4.4).
-Table 1: This table presents the sizes of the training, validation, and test sets for the datasets CIFAR-10, CIFAR-100, STL-10, OxfordIIITPet, and Flower-102.
-It also provides the number of categories (Classes) and the average number of images per category for each dataset (Per class).
-Dataset Cifar-10 Cifar-100 STL-10 OxfordIIITPet Flower-102
-Classes 10 100 10 37 102
-Train set 47500 47500 4750 3,515 969
-Val set 2500 47500 250 185 51
-Test set 10000 10000 800 3669 6149
-Per class 4750 4750 475 95 9.5
-4.1 Dataset
-In pursuit of a comprehensive understanding of the efficacy and practicality of parallel models, we meticulously selected several widely-acknowledged and representative image classification datasets for an integrated comparative analysis.
-These datasets encompass a diverse array of visual content and span various levels of difficulty, thereby furnishing us with a fertile ground to probe the performance boundaries of parallel models.
-Specifically, our study centers on the following pivotal datasets:
-1.
-CIFAR-10 & CIFAR-100 (Krizhevsky 2009): These were constructed by Alex Krizhevsky et al.
-CIFAR-10 encompasses 60,000 32x32 pixel color images across 10 classes, while CIFAR-100 extends this to 100 categories.
-2. STL-10 (Coates, Ng, and Lee 2011): STL-10 comprises 5,000 training and 8,000 testing images distributed among 10 classes, each at a 96x96 pixel resolution.
-3. Flowers-102 (Nilsback and Zisserman 2008): Focused on floral categorization, this dataset houses 102 distinct flower species, totalling 8,189 images with a per-species count ranging from 40 to 258.
+Specifically, a block in Figure.
+1 can contain multiple serial Transformer modules (it has been proven in UAT2LLMs that the mathematical form of multi-layer Transformers is also consistent with UAT), while multiple blocks operate in parallel, as illustrated in Figure.
 4.
-Oxford-IIIT Pets (Parkhi et al.
-2012): Tailored for the recognition of pet breeds (cats and dogs), this dataset contains over 7,300 images across 37 categories, characterized by varied animal poses, backgrounds, and lighting conditions.
-In Table 1, we provide more details for these dataset.
-Through experimentation on these meticulously chosen datasets, we aspire to thoroughly validate the feasibility and strengths of parallel models across a spectrum of scales, complexities, and domains in image classification tasks.
-Simultaneously, we seek to elucidate potential limitations and chart avenues for future research endeavors.
-Table 2: This table shows the predictions of Para-former-1-1, Para-former-1-6, Para-former-1-12, Para-former-1-24 on the datasets CIFAR-10, CIFAR-100, STL-10, OxfordIIITPet, and Flower-102.
-We simply represent them as Para-F-1-1, Para-F-1-6, Para-F-1-12 and Para-F-1-24, simply.
-Dataset Para-F-1-1 Para-F-1-6 Para-F-1-12 Para-F-1-24
+The number of Transformer modules within a single block is referred to as the depth, while the total number of Transformer modules across all blocks is referred to as the network’s layer count.
+$\mathbf{x}^{\prime}_{2}$ $=(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,1})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ (9)
+$+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,2})$
+Both design strategies possess the capability to dynamically adjust model parameters based on the input.
+In Figure 3, the primary mechanism is the dynamic adjustment of MHA weight parameters based on the input.
+In contrast, Figure 4 not only leverages the dynamic adjustment of MHA weight parameters but also utilizes UAT to dynamically approximate some bias parameters.
+For instance, in Eq.
+(9), $\mathbf{b}^{\prime}_{2,2}=(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$, can be seen as a two-layer UAT.
+Refer to Appendix B to see which parameters are influenced by the input.
+So, what are advantages of these strategies?
+We will talk about these in Section 4.3.
+Para-F-1-1 Para-F-1-6 Para-F-1-12 Para-F-1-24
 Cifar-10 55.95 64.90 66.88 71.53
@@
 Flower-102 27.33 32.65 37.42 35.84
-4.2 The Feasibility of Para-Former
-To validate the feasibility and characteristics of Para-Former, we conducted training and validation on several commonly used image classification datasets, with all models trained for 500 epochs.
-The experimental design involved preprocessing inputs using the common ViT techniques of patch extraction and positional encoding.
-We then used Transformer as the basic model architecture, setting all network depths to 1 and configuring Para-Former from single-layer to multi-layer setups, as shown in Figure 3.a.
-A single-layer setup corresponds to a ViT with a depth of 1.
-The results of the models are presented in Table 2.
-It is evident from Table 2 that as the number of parallel network layers increases, the model accuracy improves progressively (Except for the result of Para-F-1-24 being worse than Para-F-1-12 on the Flower-102 dataset, this variation is within a reasonable range.).
-According to UAT theory, increasing the network layers is equivalent to increasing $N$ in Equation (1), thereby enhancing the approximation capability of UAT.
-The results in Table 2 are in complete agreement with UAT theory.
-Therefore, we believe that this parallel network approach is entirely feasible.
-However, the overall prediction accuracy in Table 2 is quite low. What could be the reasons for this? We believe there are three main factors affecting the model’s predictive performance:
-1.
-Model’s Fitting Capability: This is influenced by the model’s size and design.
-Specifically, the value of $N$ in UAT and the degrees of freedom in the parameters.
-The larger the $N$ value, the stronger the UAT’s fitting capability.
-Parameter freedom refers to the number of parameters affected by the input in UAT, such as a freedom of 1 in Eq.
-8 and a freedom of 6 in Eq.
-9.
-The degrees of freedom correspond to the model’s ability to dynamically fit based on input.
-Merely increasing $N$ enhances the model’s ability to fit specific functions, but images are often diverse.
-Therefore, the network must have the ability to dynamically fit the corresponding function based on the input.
-Higher degrees of freedom mean the model can fit a wide range of features with significant differences.
-Additionally, the bias term in the network can also dynamically fit based on input using UAT, which we refer to as bias freedom.
-Models with bias freedom have a stronger dynamic fitting capability.
-The mathematical model generated by the residual network is an example of UAT with bias freedom, achieving the fitting capability of a multi-layer network with just shallow layers.
-We will talk about this problem in Section 4.3.
-2.
-Data Influence: Data is another critical factor in deep learning training.
-Since natural images are usually diverse, limited data can lead to overfitting.
-For example, in a cat-dog classification task, if dogs in the training set mostly appear against backgrounds of blue skies and grass, these backgrounds might be mistakenly considered part of the dog’s features, a phenomenon known as feature coupling.
-If cats rarely appear in such backgrounds in the training set but do in the test set, they might be misclassified as dogs.
-Training on a large, diverse dataset means that common backgrounds frequently appear with various objects or animals, such as cats, dogs, and elephants against blue skies, white clouds, houses, and trees.
-Thus, even when classifying objects, the model inherently learns background information.
-In this context, fine-tuning on the cat-dog dataset yields better results because the background does not interfere with dog feature recognition.
-Without sufficient data, even a network with strong fitting capabilities cannot perform well and is prone to overfitting.
-This problem will be considered in Section 4.4.
-Considering these factors, it is clear that simply improving the model architecture is not enough. A robust training dataset is equally important for achieving high predictive accuracy.
-Table 3: This table shows the predictions of Para-former-6-6, Para-former-2-24, Para-former-3-24, Para-former-6-24 on the datasets CIFAR-10, CIFAR-100, STL-10, OxfordIIITPet, and Flower-102.
-Dataset Para-F-6-6 Para-F-2-24 Para-F-3-24 Para-F-6-24
+Table 2: This table shows the predictions of Para-former-1-1, Para-former-1-6, Para-former-1-12, Para-former-1-24 on the datasets CIFAR-10, CIFAR-100, STL-10, OxfordIIITPet, and Flower-102.
+We simply represent them as Para-F-1-1, Para-F-1-6, Para-F-1-12 and Para-F-1-24.
+3.3 Speed-up Ratio
+We have proposed a network parallelization scheme, and in this section, we will compare the running speed of this parallelized network with that of traditional serial networks.
+It is important to note that due to differences in hardware and implementation methods, providing an objective comparison through experiments can be challenging.
+Therefore, we will focus on a theoretical analysis.
+Based on the network structure shown in Figure 3 and 4, the inference time of Para-Former is only affected by its network depth.
+Assume we use a Para-Former with a depth of $M$ and a layer count of $L$, compared to a traditional Transformer network with a depth of $N$ layers, where $L>>N$.
+In this case, the running speed of the parallel network would be $N/M$ times faster than that of the serial network.
+This is because the speed of the parallel network is solely related to the depth of the Para-Former.
+As the serial network depth increases, the acceleration effect becomes more pronounced.
+4 Experiments
+Our experimental approach is to first verify the feasibility of the parallel network theory and its adherence to the UAT approximation properties (The larger the value of $N$, the more powerful the model becomes, resulting in better classification performance.) using several commonly used public image classification datasets (Section 4.2).
+We then explore the impact of network depth on the Para-Former performance (Section 4.3), followed by an investigation into the influence of the dataset (Section 4.4).
+We conducted experiments with various Para-Former models, where the naming convention uses the first number to indicate the depth and the second number to indicate the number of layers.
+For instance, Para-former-1-12 denotes a model with 1 depth and 12 layers.
+4.1 Dataset
+In pursuit of a comprehensive understanding of the efficacy and practicality of parallel models, we meticulously selected several widely-acknowledged and representative image classification datasets for an integrated comparative analysis.
+These datasets encompass a diverse array of visual content and span various levels of difficulty, thereby furnishing us with a fertile ground to probe the performance boundaries of parallel models.
+Specifically, our study centers on the following pivotal datasets: CIFAR-10, CIFAR-100 (Krizhevsky 2009), STL-10 (Coates, Ng, and Lee 2011), Flowers-102 (Nilsback and Zisserman 2008) and Oxford-IIIT Pets (Parkhi et al.
+2012).
+In Table 1, we provide more details for these datasets.
+Para-F-6-6 Para-F-2-24 Para-F-3-24 Para-F-6-24
 Cifar-10 79.39 74.75 77.21 79.92
@@
 Flower-102 37.97  39.17 40.12 40.51
-4.3 Model Depth Effect
-Based on the conclusions from Section 4.2, we will explore the impact of network depth on predictive performance in this section.
-According to Eq.
-8 and Eq.
-9, we know that model depth primarily affects the degrees of freedom of the model parameters in UAT.
-Therefore, we designed Para-Former models with 24 layers and varying depths of 2, 3, and 6, and compared their predictive performance with that of a ViT with a depth of 6.
-The results are shown in Table 3.
-It is evident that all results surpass those of Para-F-1-24 in Table 2 because the networks in Table 3 have higher degrees of freedom in their parameters.
-Notably, Para-F-6-6 consistently outperforms Para-F-2-24 and Para-F-3-24, indicating that the higher the degrees of freedom of the parameters, the stronger the model’s performance.
-However, the model cannot be too deep, so we must balance network depth and layer count.
-Another issue is that almost all models hit a performance bottleneck.
-Particularly, the prediction accuracy for OxfordIIITPet and Flower-102 is very poor.
-This is because these datasets are relatively small and exhibit high data diversity.
+Table 3: This table shows the predictions of Para-former-6-6, Para-former-2-24, Para-former-3-24, Para-former-6-24 on the datasets CIFAR-10, CIFAR-100, STL-10, OxfordIIITPet, and Flower-102.
+4.2 The Feasibility of Para-Former
+To validate the feasibility and characteristics of Para-Former, we conducted training and validation on several commonly used image classification datasets, with all models trained for 500 epochs.
+The experimental design involved preprocessing inputs using the common ViT techniques of patch extraction and positional encoding.
+We then used Transformer as the basic model architecture, setting all network depths to 1 and configuring Para-Former from single-layer to multi-layer setups, as shown in Figure 3 (We evaluated the models’ performance using the accuracy metric, calculated as $\text{accuracy}=100\times\frac{\text{correct}}{\text{total}}$).
+A single-layer setup corresponds to a ViT with a depth of 1.
+The results of the models are presented in Table 2.
+It is evident from Table 2 that as the number of parallel network layers increases, the model accuracy improves progressively (Except for the result of Para-F-1-24 being worse than Para-F-1-12 on the Flower-102 dataset, this variation is within a reasonable range.).
+According to UAT theory, increasing the network layers is equivalent to increasing $N$ in Eq.
+(1), thereby enhancing the approximation capability of UAT.
+The results in Table 2 are in complete agreement with UAT theory.
+Therefore, we believe that this parallel network approach is entirely feasible.
+However, the overall prediction accuracy in Table 2 is quite low. What could be the reasons for this? We believe there are two main factors affecting the model’s predictive performance:
+1.
+Model Fitting Capability: This is influenced by the model’s size and design, particularly the value of $N$ in the UAT and the degrees of freedom of the parameters.
+A larger $N$ enhances the UAT’s ability to fit functions.
+The degrees of freedom refer to the number of parameters in the UAT that are influenced by the input.
+For instance, the degrees of freedom in Equation (8) is 2, while in Equation (9), it is 6 (see Appendix B).
+The degrees of freedom represent the model’s ability to dynamically fit corresponding functions based on the input.
+Simply increasing $N$ can improve the model’s ability to fit specific functions, but images are usually diverse.
+Thus, the network must have the capability to dynamically fit different functions based on the input.
+Higher degrees of freedom allow the model to fit multiple different functions according to the input.
+Additionally, the degrees of freedom vary within the Transformer: they mainly consist of weights and biases.
+The degrees of freedom for weights stem from the MHA, primarily putting linear transformations based on the input.
+The degrees of freedom for biases come from residual connections, which can dynamically fit the bias terms using UAT.
+The deeper the residual layers, the more UAT layers correspond to the bias terms, leading to stronger dynamic fitting capabilities.
+As a result, a shallow serial network can achieve the fitting ability of a multi-layer parallel network with a depth of 1.
+We will explore this issue further in Section 4.3.
+2.
+Data Influence: Data is another crucial factor in deep learning training.
+Since natural images are typically diverse, limited data can lead to overfitting.
+For example, in a cat-dog classification task, if dogs in the training set mostly appear against backgrounds of blue skies and grass, these backgrounds might be mistakenly considered features of dogs.
+If cats rarely appear in such backgrounds in the training set but do in the test set, they might be misclassified as dogs.
+Training on a large, diverse dataset means common backgrounds frequently appear with various objects or animals, such as cats, dogs, and elephants alongside blue skies, white clouds, houses, and trees.
+Therefore the model will learn background information unconsciously when doing classification tasks.
+Then fine-tuning on a smaller cat-dog dataset will yield better results because the background won’t interfere with the recognition of dog features.
+Without sufficient data, even a network with strong fitting capabilities cannot perform well and is prone to overfitting.
+We will consider this issue in Section 4.4.
 Figure 5: OxfordIIITPet dataset’s accuracy on different classes.
 Figure 6: Flower-102 dataset’s accuracy on different classes.
+4.3 Model Depth Effect
+Based on the conclusions from Section 4.2, this section explores the impact of network depth on prediction performance.
+According to equations (8) and (9), the depth of the model primarily affects the degrees of freedom of the model parameters within the UAT.
+Therefore, we designed different versions of the Para-Former model with 24 layers and depths of 2, 3, and 6, comparing their predictive performance against a ViT with a depth of 6.
+The results are shown in Table 3.
+The degrees of freedom for Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24 are 48, 18, 72, 72, and 72, respectively (see Appendix B).
+Why does Para-F-6-6, despite having the lowest degrees of freedom, only perform worse than Para-F-6-24?
+This is because the degrees of freedom encompass both the freedom of weights and biases.
+The freedom of weights can be understood as linear transformations derived from inputs, whereas the freedom of biases results from the approximation by multiple layers of UAT, which allows for stronger dynamic fitting capabilities.
+The deeper the layers, the stronger the approximation ability.
+Thus, Para-F-6-24 performs the best, followed by Para-F-6-6, as the number of UAT layer of their bias terms is largest (see Appendix B).
+In summary, deeper models have stronger approximation capabilities, but there is a limit to how deep a model can go, necessitating a balance between network depth and layer count.
 4.4 Dataset Effect
-Table 4: This table shows the results of fine-tuning using Google’s official pre-trained parameters. The pre-trained parameters were obtained from training the ViT model on the ImageNet dataset.
-Dataset Cifar-10 Cifar-100 STL-10 OxfordIIITPet Flower-102
+Cifar-10 Cifar-100 STL-10 OxfordIIITPet Flower-102
 Para-F-12-12 98.62 91.36 99.47 91.87 96.89
-Current deep learning models generally perform well on the aforementioned datasets.
-However, as shown in Table 3, the prediction accuracy of Para-Former is not very high, especially on the OxfordIIITPet and Flower-102 datasets.
+Table 4: This table shows the results of fine-tuning using Google’s official pre-trained parameters of ViT.
+The pre-trained parameters were obtained from training the ViT model on the ImageNet dataset.
+Figure 7: The fine-tuning process from the perspective of UAT.
+After increasing the depth of the Para-Former, the results have improved to some extent.
+However, as shown in Table 3, almost all models encounter a performance bottleneck, especially on the OxfordIIITPet and Flower-102 datasets.
 What is the fundamental reason for this poor performance?
-In this section, we will explain the relationship between deep learning and data.
-First, the results were obtained using the validated ViT model (Para-F-1-1 and Para-Former-6-6), which performed similarly to Para-Former.
-Given that the ViT model has been extensively validated and is a commonly used model architecture, the effectiveness of the models is unquestionable.
-So, why is the performance so poor?
-The root cause lies in the data itself: the model’s performance is generally directly proportional to the amount of available data.
-Take the Flower-102 dataset, for example, which includes 102 classes of flowers but has only 1,020 images in the training set, with some reserved for validation.
-This is far from sufficient to train deep learning models.
-Why, then, does the OxfordIIITPet dataset, which has more data per class on average than Flower-102, perform even worse?
-This is due to the diversity of the data (i.e., the similarity of features within the same class) and the variety of backgrounds.
-Figures 4 and 5 show some results from the OxfordIIITPet and Flower-102 datasets.
-Both figures indicate that prediction results are worse for categories with diverse backgrounds or inherent diversity, as seen in Figures 4.b and 5.b.
-Conversely, results are relatively better when the background is consistent but the species vary, as shown in Figures 4.a and 5.a.
-The best results are achieved when both the background and species variation are minimal, as shown in Figures 4.c and 5.c.
-Although the Flower-102 dataset has fewer images, its backgrounds are more consistent and there is less variation among the species.
-In contrast, animals in the OxfordIIITPet dataset can appear in various forms, while flowers in the Flower-102 dataset are usually photographed from relatively fixed angles, leading to better performance.
+This is likely because these datasets are relatively small and exhibit high data diversity.
+In this section, we will solve this problem.
+First, the experimental results are based on validated ViT models (Para-F-1-1 and Para-Former-6-6 are equal to ViT with 1 and 6 depth.), whose performance is similar to that of the Para-Former.
+Given that ViT models are widely validated and are among the commonly used model architectures, their effectiveness is beyond doubt.
+So why is the performance so poor?
+The fundamental reason lies in the data itself: the model’s performance is typically directly related to the amount of available data.
+Take the Flower-102 dataset as an example, which contains 102 classes of flowers but only 1,020 training images, with some reserved for validation.
+This is far from sufficient for training deep learning models.
+Why is the performance even worse on the OxfordIIITPet dataset, which has a higher average number of images per class?
+This is due to the diversity of the data (i.e., the similarity of features within the same class) and the variability of the background.
+Figures 5 and 6 show some results from the OxfordIIITPet and Flower-102 datasets.
+Both figures indicate that for categories with diverse backgrounds and inherent feature diversity, the prediction results are poor, as shown in Figures 5.b and 6.b.
+In contrast, when the background is consistent but the object morphology is diverse, the results are relatively better, as shown in Figures 5.a and 6.a.
+The best results are achieved when both the background and object morphology are minimal, as shown in Figures 5.c and 6.c.
+Although the Flower-102 dataset has fewer images, its backgrounds are more consistent, and the differences between species are smaller.
+Conversely, animals in the OxfordIIITPet dataset can appear in various forms, whereas flowers in the Flower-102 dataset are usually photographed from relatively fixed angles, leading to better performance.
 Does this mean that deep learning models are not powerful enough or lack generalization and reasoning capabilities?
-On the contrary, achieving around 25% accuracy on OxfordIIITPet and about 40% on Flower-102 with such limited data actually demonstrates the strength and generalization capability of deep learning models.
-A common misconception about deep learning is that humans, even if unfamiliar with certain species, can accurately distinguish OxfordIIITPet and Flower-102 images by carefully examining them, thus believing that deep learning lacks generalization or human-like inferencing ability.
-This viewpoint is flawed.
-Assuming the human brain functions as a UAT model, it can automatically disregard background information before seeing these images, as this background is frequently encountered in daily life, and the UAT in the human brain has already fitted this information.
-Therefore, humans only need to focus on specific information and then use it to train their brains to find similarities and differences.
-For instance, a 10-year-old child’s brain, observing the world at 60 frames per second, would have processed about $1\times 10^{10}$ high-resolution images by the age of 10, many of which include common objects and backgrounds in daily life.
-Therefore, criticizing deep learning models for being unable to train on small datasets and lacking universality and generalization is unfair.
-Thus, we should train on large datasets and then fine-tune on smaller ones.
-This is also the current deep learning training method, showcasing the powerful generalization capability of deep learning models, especially when there is some correlation between datasets.
-Following this fine-tuning approach, we fine-tuned the pre-trained ViT model on the aforementioned datasets for just 10 epochs.
-The results, as shown in Table 4, indicate a significant improvement in prediction performance.
-Figure 6 shows the rationality behind fine-tuning.
-Since we essentially adjust parameters based on the input to produce target results using UAT, if there is a correlation in the data, the corresponding parameters derived from the input ($\mathbf{\alpha}_{i},\mathbf{W}_{i}$, and $\mathbf{\theta}_{i};i=1,\cdot,N$) will be similar.
-Therefore, fine-tuning is feasible.
-Figure 6: The fine-tuning process from the perspective of UAT.
+Quite the opposite.
+With such limited data, deep learning models achieved about 25% accuracy on OxfordIIITPet and about 40% on Flower-102, actually demonstrating the strength and generalization ability of deep learning models.
+A common misconception about deep learning is that even if people are unfamiliar with certain species, by careful examination, they can accurately distinguish between OxfordIIITPet and Flower-102 images, leading to the belief that deep learning lacks generalization or human-like reasoning capabilities.
+This view is flawed.
+Suppose the human brain operates like a UAT model; it can automatically ignore background information before seeing these images because such backgrounds are frequently encountered in daily life, and the UAT in the brain has already fitted this information.
+Therefore, people only need to focus on specific information and then use this information to train the brain to identify similarities and differences.
+For example, a 10-year-old child’s brain, observing the world at 10 frames per second (Read and Meyer 2000), would have processed over $1.5\times 10^{9}$ high-resolution images by age 10 (Lu et al.
+2017), many containing common objects and backgrounds from daily life.
+Hence, it is unfair to criticize deep learning models for being unable to train on small datasets and for lacking universality and generalization ability.
+Therefore, we should train on large datasets and then fine-tune on smaller ones.
+This is also the current method for deep learning training, demonstrating the powerful generalization capabilities of deep learning models, especially when there is some correlation between the datasets.
+Following this approach, we fine-tuned a pre-trained ViT model using only 10 epochs.
+The results, as shown in Table 4, demonstrate significant improvements in prediction performance.
+Figure 7 illustrates the rationale for fine-tuning.
+The goal, whether in training or prediction, is to obtain the target results based on the input.
+This can be understood as function fitting, expressed as $\mathbf{y}=\mathbf{G}(\mathbf{x}|\mathbf{A},\mathbf{W},\mathbf{\Theta})$, where $\mathbf{A}={\mathbf{\alpha}_{1}...\mathbf{\alpha}_{N}}$, $\mathbf{W}={\mathbf{W}_{1}...\mathbf{W}_{N}}$, and $\mathbf{\Theta}={\mathbf{\theta}_{1}...\mathbf{\theta}_{N}}$.
+The purpose of training is to determine $\mathbf{A},\mathbf{W},\mathbf{\Theta}$, which can dynamically change based on the input.
+This means that if the background is similar but the objects are different, the resulting parameters will differ.
+If the model can make correct predictions, it implies that the model can distinguish between the background and the animals.
+If it could not be distinguished, it might incorrectly use the background as a feature of the animals, leading to the same classification.
+Moreover, if the model can correctly classify the animals, it indicates that the model can identify the features of the animals.
+When the backgrounds are different, because natural backgrounds are generally similar, extensive training can differentiate background information.
+Thus, if the model can make right predictions, it suggests that the model has comprehensively grasped the animal features, even with significant environmental variations.
+Training on ImageNet follows this process.
+A new dataset is just a new series of combinations of various similar backgrounds and animals which have been seen in ImageNet.
+If the model has already learned sufficient background and animal information on ImageNet, it can dynamically fit the results based on the input.
+Therefore, only minor adjustments specific to new data are required during fine-tuning.
 5 Conclusion
-In this paper, we propose a parallel network based on the Universal Approximation Theorem (UAT) and validate its effectiveness on large datasets.
+In this paper, we propose a parallel network based on the UAT and validate its effectiveness on large datasets.
 We analyze the characteristics of both serial and parallel networks.
-Serial networks, especially those based on residual connections, often have higher degrees of parameter freedom, leading to stronger generalization capabilities.
+Serial networks based on residual connections, often have higher degrees of parameter freedom, leading to stronger generalization capabilities.
 However, as the network depth increases, the inference time also becomes longer.
@@
 Additionally, we analyze the impact and reasons data have on deep learning.
-From the UAT perspective, training on large datasets essentially enhances the model’s ability to fit data.
-Appendix A The UAT Format of Transformer
-A.1 The UAT Format of Transformer-based ViTs
-Eq.(1) and (2) give an example of the matrix-vector format of MHA and FFN.
-Eq.
+From the UAT perspective, training on large datasets essentially enhances the model’s ability to fit diverse data.
+Appendix A The UAT Format of Transformer-based ViTs
+In this section, we will demonstrate that the mathematical form of single-layer and two-layer Transformers is equivalent to the UAT.
+Firstly, the basic form of a Transformer is shown in Figure 2 of main text.
+We represent the MHA and FFN in the Transformer in matrix-vector form as Eq.
 (10) and Eq.
-(4) give the corresponding UAT format of $\mathbf{x}_{i+1}$ and $\mathbf{x}_{i+2}$.
-It is obvious that the bias and weight terms in the UAT corresponding to Transformed can change dynamically based on the input.
-Due to the complexity of the variables involved, we will combine some terms to align the resulting equations with the standard UAT form.
-To distinguish these parameter variables, we will keep the original variables unchanged.
-If a parameter variable is influenced only by other variables, we will denote it with a bar on top, such as $\bar{\mathbf{W}}$.
-If it is influenced by the input $\mathbf{x}_{i}$, we will mark it with a hat, such as $\hat{\mathbf{W}}$.
-$MHA(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}$ (1)
-$FFN(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}\$ (2)
-$\mathbf{x}^{\prime}_{i+1}$ $=\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma[\mathbf{W}^{\prime}_{i,2}(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i})+\mathbf{b}^{\prime}_{i,2}]+\mathbf{b}^{\prime}_{i,3}$ (3)
-$=\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma(\underline{\mathbf{W}^{\prime}_{i,2}\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}$
-$=\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}$
-$\mathbf{x}^{\prime}_{i+2}$ $=\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i+1}+\mathbf{W}^{\prime}_{i+1,3}\sigma(\hat{\mathbf{W}^{\prime}_{i+1,1}}\mathbf{x}^{\prime}_{i+1}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{b}^{\prime}_{i+1,3}$ (4)
-$=\mathbf{W}^{\prime}_{i+1,1}[\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}]$
-$+\mathbf{W}^{\prime}_{i+1,3}\sigma\{\hat{\mathbf{W}^{\prime}_{i+1,1}}[\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}]+\mathbf{b}^{\prime}_{i+1,2}\}+\mathbf{b}^{\prime}_{i+1,3}$
-$=\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}^{\prime}_{i,3}}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{b}^{\prime}_{i,3}}$
-$+\mathbf{W}^{\prime}_{i+1,3}\sigma\{\underline{\hat{\mathbf{W}^{\prime}_{i+1,1}}\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\underline{\hat{\mathbf{W}^{\prime}_{i+1,1}}\mathbf{W}^{\prime}_{i,3}}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\underline{\hat{\mathbf{W}^{\prime}_{i+1,1}}\mathbf{b}^{\prime}_{i,3}+\mathbf{b}^{\prime}_{i+1,2}}\}+\mathbf{b}^{\prime}_{i+1,3}$
-$=\hat{\mathbf{W}^{\prime}_{i+1,2}}\mathbf{x}^{\prime}_{i}+\hat{\mathbf{W}^{\prime}_{i+1,3}}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{W}^{\prime}_{i+1,3}\sigma\{\hat{\mathbf{W}^{\prime}_{i+1,4}}\mathbf{x}^{\prime}_{i}+\underline{\hat{\mathbf{W}^{\prime}_{i+1,5}}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\hat{\mathbf{b}^{\prime}_{i+1,2}}}\}+\underline{\hat{\mathbf{b}^{\prime}_{i,3}}+\mathbf{b}^{\prime}_{i+1,3}}$
-$=\hat{\mathbf{W}^{\prime}_{i+1,2}}\mathbf{x}^{\prime}_{i}+\hat{\mathbf{W}^{\prime}_{i+1,3}}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{W}^{\prime}_{i+1,3}\sigma\{\hat{\mathbf{W}^{\prime}_{i+1,4}}\mathbf{x}^{\prime}_{i}+\hat{\mathbf{b}^{\prime}_{i+1,4}}\}+\hat{\mathbf{b}^{\prime}_{i+1,5}}$
+(11), respectively (for specific details, refer to UAT2LLMs).
+To distinguish between the operations and variables in the Transformer and those in the matrix-vector form, we add a prime () symbol to the operations and variables in the matrix-vector form.
+It is important to emphasize that the parameters of the MHA are influenced by the input.
+$MHA^{\prime}(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}$ (10)
+$FFN^{\prime}(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{b}^{\prime}_{i+1,3}\$ (11)
+Next, the result of a single-layer Transformer $\mathbf{x}^{\prime}_{1}$ is:
+$\mathbf{x}^{\prime}_{1}$ $=\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{W}^{\prime}_{1,3}\sigma[\mathbf{W}^{\prime}_{1,2}(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0})+\mathbf{b}^{\prime}_{1,2}]+\mathbf{b}^{\prime}_{1,3}$ (12)
+$=\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{W}^{\prime}_{1,3}\sigma(\underline{\mathbf{W}^{\prime}_{1,2}\mathbf{W}^{\prime}_{1,1}}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})+\mathbf{b}^{\prime}_{1,3}$
+Let $\mathbf{W}^{\prime}_{1,2}=\underline{\mathbf{W}^{\prime}_{1,2}\mathbf{W}^{\prime}_{1,1}}$, so we obtain:
+$\mathbf{x}^{\prime}_{1}=(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ (13)
+where $\mathbf{W}^{\prime}_{1,1}$ and $\mathbf{W}^{\prime}_{1,2}$ are parameters influenced by the input, thus having a degree of freedom of 2.
+$\mathbf{x}^{\prime}_{2}$ $=\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{1}+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{2,2})+\mathbf{b}^{\prime}_{2,3}$ (14)
+$=(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{2,3})+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{2,2})$
+=$=\{\mathbf{W}^{\prime}_{2,1}[(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})]+\mathbf{b}^{\prime}_{2,3}\}$
+$+\mathbf{W}^{\prime}_{2,3}\sigma\{\mathbf{W}^{\prime}_{2,2}[(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})]+\mathbf{b}^{\prime}_{2,2}\}$
+=$=\{\mathbf{W}^{\prime}_{2,1}(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{b}^{\prime}_{2,3}+\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})\}$
+$+\mathbf{W}^{\prime}_{2,3}\sigma\{\mathbf{W}^{\prime}_{2,2}({\mathbf{W}}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{b}^{\prime}_{2,2}+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})\}$
+=$=\{(\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,1}}\mathbf{x}^{\prime}_{0}+\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,3}})+\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})\}$
+$+\mathbf{W}^{\prime}_{2,3}\sigma\{\underline{\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}}\mathbf{x}^{\prime}_{0}+\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}\}$
+In Eq.
+(14), we provide the matrix-vector format of 2-layer Transformer.
+Let $\mathbf{W}^{\prime}_{2,1}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,1}}$, $\mathbf{b}^{\prime}_{2,1}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,3}}$, $\mathbf{W}^{\prime}_{1,3}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}}$, $\mathbf{W}^{\prime}_{2,2}=\underline{\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}}$, and $\mathbf{b}^{\prime}_{2,2}=\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}$.
+Thus, Eq.
+(14) can be written as:
+$\mathbf{x}^{\prime}_{2}$ $=(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,1})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ (15)
+$+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,2})$
+where $\mathbf{W}^{\prime}_{2,1}$,$\mathbf{b}^{\prime}_{2,1}$,$\mathbf{W}^{\prime}_{1,2}$,$\mathbf{W}^{\prime}_{1,3}$,$\mathbf{W}^{\prime}_{2,3}$, and $\mathbf{b}^{\prime}_{2,2}$ are influenced by the input, so the degree of freedom is 6.
+Another issue is that in $\mathbf{x}^{\prime}_{1}$ and $\mathbf{x}^{\prime}_{2}$, there is a term $(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,3})$ and $(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,1})$ which doesn’t generally conform to the standard terms in UAT.
+We can consider this term as a special case and regard it as a layer in UAT.
+Firstly, when the number of network layers increases, it effectively increases $N$ in UAT.
+Given UAT’s strong approximation capabilities, adding multiple instances of this term can be viewed as a linear transformation, thus it doesn’t affect the overall approximation power of UAT.
+Furthermore, this term can be considered a special form in UAT.
+Taking ReLU as an example activation function, $(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,3})$ can be expressed as $\text{ReLU}(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,3})-\text{ReLU}(-(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,3}))$.
+In summary, we have shown that the mathematical forms of one-layer and two-layer Transformers can be seen as 2 and three-layer UAT.
+Appendix B The Degree of Freedom of Different Layers Transformer
+According to UAT2LLMs, the UAT format of multi-layer Transformer can be written as:
+$\mathbf{x}_{i+1}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,1})+\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$ (16)
+where $\mathbf{b}^{\prime}_{j,2}=(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})+\mathbf{W}^{\prime}_{j,2}UAT_{j-1}^{R}$, where $UAT_{j-1}^{R}=\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,3}\sigma(\mathbf{W}^{\prime}_{k,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,2})$.
+The term $\mathbf{b}^{\prime}_{j,2}$ is approximated by a $j$-layer of UAT with $\mathbf{x}_{0}$ as input.
+All $\mathbf{W}^{\prime}_{i,1}$, $\mathbf{W}^{\prime}_{j,2}$, and $\mathbf{W}^{\prime}_{j,3}$ parameters in layer $i$, where $j=1,\ldots,i$, are dynamically adjusted based on the input.
+Because they are affected by the weight of MHA in each layer.
+Why does the MHA operation affect the weight parameters in a linear manner?
+According to Eq.
+(12) and Eq.
+(13), this is because $\mathbf{W}^{\prime}_{1,2}=\underline{\mathbf{W}^{\prime}_{1,2}\mathbf{W}^{\prime}_{1,1}}$, where $\mathbf{W}^{\prime}_{1,1}$ represents the MHA parameters, which are influenced by the input, while $\mathbf{W}^{\prime}_{1,2}$ is a defined parameter.
+As a result, the parameter $\mathbf{W}^{\prime}_{1,2}$ in Eq.
+(13) becomes input-dependent, and this influence occurs through a linear transformation.
+Similarly, in a multi-layer Transformer, multiple layers of weight parameters are derived through similar transformations.
+Therefore, it is easy to deduce that the degrees of freedom for Transformers with 1, 2, 3, and 6 layers are 2, 6, 9, and 18, respectively.
+For Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24, the degrees of freedom are shown in Table 1.
+At first glance, it seems that the degrees of freedom for Para-F-1-24, Para-F-2-24, Para-F-3-24, and Para-F-6-24 all exceed that of Para-F-6-6.
+Why then does only Para-F-6-24 consistently outperform Para-F-6-6?
+Para-F-1-24 Para-F-6-6 Para-F-2-24 Para-F-3-24 Para-F-6-24
+Degree of Freedom 2*24 18*1  6*12 9*8 18*4
+Table 1: The degree of freedom of Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, Para-F-6-24.
+We use the the degree of freedom of serial Transformer with the depth 1, 6, 2, and 3 to multiply the number of serial Transformer.
+The reason lies in the nature of the degrees of freedom.
+Taking $\mathbf{W}^{\prime}_{2,2}$ and $\mathbf{b}^{\prime}_{2,2}$ from Eq.
+(15) as examples, $\mathbf{W}^{\prime}_{2,2}=\underline{\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}}$, where $\mathbf{W}^{\prime}_{1,1}$ is an input-dependent parameter from MHA, and $\mathbf{W}^{\prime}_{2,2}$ is a defined parameter, becoming fixed post-training.
+Thus, the degrees of freedom for $\mathbf{W}^{\prime}_{2,2}$ in Eq.
+(15) are imparted through a linear transformation with the MHA parameter $\mathbf{W}^{\prime}_{1,1}$.
+Additionally, $\mathbf{b}^{\prime}_{2,2}=\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}$.
+Hence, $\mathbf{b}^{\prime}_{2,2}$ in Eq.
+(15) can be understood as being approximated by a two-layer UAT, showcasing stronger dynamic transformation capabilities based on inputs.
+Due to UAT’s powerful capacity, $\mathbf{b}^{\prime}_{2,2}$ can dynamically fit specific functions based on inputs.
+In a single-layer Transformer, as seen in Eq.
+(12), $\mathbf{b}^{\prime}_{1,3}$ and $\mathbf{b}^{\prime}_{1,2}$ are defined parameters, so the number of bias terms approximated using UAT is zero.
+For a two-layer Transformer, as shown in Eq.
+(15), the number of bias terms approximated using UAT is two, specifically $\mathbf{b}^{\prime}_{2,1}$ and $\mathbf{b}^{\prime}_{2,2}$, with the UAT layers approximating these bias terms being 1 and 2, respectively.
+According to Eq.(16), in a Transformer with $i+1$ layers, $\mathbf{b}_{i+1,1}$ is approximated by a single-layer UAT, while $\mathbf{b}^{\prime}_{j,2}$ is approximated by a UAT of $j$ layers for $j>1$.
+It is easy to deduce that for a three-layer Transformer, the number of bias terms approximated using UAT is three, with the UAT layers being 1, 2, and 3.
+For a six-layer Transformer, the number of bias terms approximated using UAT is six, with the UAT layers being 1, 2, 3, 4, 5, and 6.
+The UAT layers of bias terms for Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24 are shown in Table 2.
+Para-F-1-24 Para-F-6-6 Para-F-2-24 Para-F-3-24 Para-F-6-24
+UAT Layers (0)*24 (1+2+3+4+5+6)*1 (1+2)*12 (1+2+3)*8 (1+2+3+4+5+6)*4
+Table 2: The UAT layer count corresponding to the bias terms in Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24 is calculated by summing the UAT layers associated with each bias term within a serial Transformer.
+Since each serial Transformer has multiple bias terms, with each term corresponding to different UAT layer counts, we represent the UAT layers by summing the UAT layers for each bias term.
+For example, ”1+2+3” indicates that three bias terms are approximated using UAT, with 1, 2, and 3 layers respectively.
+Additionally, since multiple serial Transformers are used in parallel, we multiply the sum by the number of parallel Transformers.
+In summary, more layers mean more degrees of freedom in bias terms, leading to stronger model fitting capabilities.
+To achieve performance comparable to that of deep serial networks while ensuring model fitting ability and fast inference, parallel networks need to strike a balance between network depth and layer count.
+Parallel networks are designed for extremely large models.
+For instance, if a serial network with 1 million layers can be replaced by a parallel network with a depth of 10 and 100 million layers, the speedup factor can reach up to 100,000 times.
+Sure, we could improve the design of the Para-former to achieve speedup computing with fewer layers.

### v5

@@
-Universal Approximation Theory: Foundations for Parallelism in Neural Networks
+Dynamic Universal Approximation Theory Foundation for Parallelism in CV Neural Networks
 Abstract
@@
 This is unacceptable if deep learning models go deeper.
-Therefore, this paper proposes a deep learning parallelization strategy based on the Universal Approximation Theorem (UAT).
+Therefore, this paper proposes a deep learning parallelization strategy based on the Dynamic Universal Approximation Theorem (DUAT).
 From this foundation, we designed a parallel network called Para-Former to test our theory.
@@
 1 Introduction
-Currently, deep learning is advancing towards larger models and datasets, driven by the success of ChatGPT (Radford and Narasimhan 2018; Brown et al.
-2020; Achiam et al.
-2023), Llama (Touvron et al.
-2023), and ViT (Dosovitskiy et al.
-2020), among others.
-These large models demonstrate significantly enhanced performance compared to their predecessors.
-According to UAT2LLMs (Wang and Li 2024b) and UAT2CVs (Wang and Li 2024b), residual-based convolutional neural networks (CNNs) and Transformer-based large language models are specific implementations of the UAT (Cybenko 2007; Hornik, Stinchcombe, and White 1989), capable of dynamically approximating functions based on inputs.
-From the perspective of UAT, larger models inherently possess stronger dynamic fitting capabilities (Wang and Li 2024b, a).
-Therefore, the trend towards larger models is undoubtedly the future direction of deep learning, and large datasets are essential for training these models to ensure their approximations accurately reflect real-world complexities.
+Currently, deep learning is advancing toward larger-scale models and datasets, a trend driven by the successes of models such as ChatGPT Radford & Narasimhan (2018); Brown et al.
+(2020); Achiam et al.
+(2023), Llama Touvron et al.
+(2023), and ViT Dosovitskiy et al.
+(2020).
+These large models exhibit significantly enhanced performance compared to their predecessors.
+According to DUAT2LLMs Wang & Li (2024b) and DUAT2CVs Wang & Li (2024a), both residual-based convolutional neural networks (CNNs) and large-scale Transformer-based models are essentially DUAT functions Wang & Li (2024b; a).
+DUAT, an evolution of the Universal Approximation Theory (UAT) Cybenko (1989); Hornik et al.
+(1989), enables dynamic function approximation based on input data.
+From the DUAT perspective, larger models inherently possess a greater dynamic fitting capacity Wang & Li (2024b; a).
+Consequently, the shift toward larger models is undoubtedly the future of deep learning, and large datasets are essential for training these models to ensure their approximations accurately reflect real-world complexity.
 However, this trend towards larger models and datasets inevitably leads to increased demands for computational resources and extended training times.
@@
 As network layers increase to tens or even hundreds of thousands, inference times can extend significantly, potentially requiring several minutes or hours for a single batch, making training large models on large datasets exceedingly difficult.
-Current solutions include model optimization techniques such as quantization (Jacob et al.
-2017; Lu, Zhong, and Yang 2022; Guo, Wang, and Cui 2021; Lin et al.
-2023), pruning (Ma, Fang, and Wang 2023; Xia et al.
-2023), and knowledge distillation (Gou et al.
-2020; Gu et al.
-2021).
-Additionally, specialized accelerators (Jouppi et al.
-2017; Elbtity, Chandarana, and Zand 2024) and edge computing (Shi et al.
-2016; Gale et al.
-2020), as well as parallel and distributed computing (Shallue et al.
-2018; Ben-Nun and Hoefler 2018; Sergeev and Balso 2018), offer potential solutions.
+Current solutions include model optimization techniques such as quantization Jacob et al.
+(2017); Lu et al.
+(2022); Guo et al.
+(2021); Lin et al.
+(2023), pruning Ma et al.
+(2023); Xia et al.
+(2023), and knowledge distillation Gou et al.
+(2020); Gu et al.
+(2021).
+Additionally, specialized accelerators Jouppi et al.
+(2017); Elbtity et al.
+(2024) and edge computing Shi et al.
+(2016); Gale et al.
+(2020), as well as parallel and distributed computing Shallue et al.
+(2018); Ben-Nun & Hoefler (2018); Sergeev & Balso (2018), offer potential solutions.
 However, these methods do not fundamentally address the inference delay caused by the increasing number of serial network layers.
@@
 This design originated from early computer vision networks, which used convolutional layers with small receptive fields to accommodate the local correlations and spatiotemporal invariance of image data.
-Achieving a larger receptive field required deeper network layers, as exemplified by architectures like AlexNet (Krizhevsky, Sutskever, and Hinton 2012) and VGG (Simonyan and Zisserman 2014).
-The introduction of ResNe t (He et al.
-2015) marked a paradigm shift, as the residual structure could dynamically fit functions based on input (Wang and Li 2024b, a), influencing subsequent network designs and establishing the tradition of multi-layer stacking.
+Achieving a larger receptive field required deeper network layers, as exemplified by architectures like AlexNet Krizhevsky et al.
+(2012) and VGG Simonyan & Zisserman (2014).
+The introduction of ResNet He et al.
+(2015) marked a paradigm shift, as the residual structure could dynamically fit functions based on input Wang & Li (2024b; a), influencing subsequent network designs and establishing the tradition of multi-layer stacking.
 While this design was manageable with fewer layers, the increasing depth and number of parameters in modern networks have created a critical issue: slow computation efficiency due to the need for sequential layer-by-layer calculations during inference.
 This computation method inevitably leads to delays, which become more severe as network depth increases, significantly constraining the progress of computer vision and natural language processing technologies.
-Therefore, it is imperative to explore parallelization techniques for deep learning networks, enabling parallel computation across multiple layers.
-This approach is key to accelerating the inference process, breaking the time bottleneck, and advancing computer vision and natural language processing technologies towards greater efficiency and speed.
-To develop such parallel computation, we must start from the fundamental theory of deep learning: the UAT.
-The network must be designed in accordance with UAT principles and must satisfy the requirements outlined in the papers UAT2LLMs (Wang and Li 2024b) and UAT2CVs (Wang and Li 2024a), which emphasize the ability to dynamically fit functions based on data.
-With this in mind and leveraging existing network structures, we designed a parallel network, Para-Former, to validate the feasibility of our theory.
+Therefore, exploring parallelization techniques for deep learning networks—enabling parallel computation across multiple layers—is crucial for accelerating inference, overcoming time bottlenecks, and advancing computer vision and natural language processing technologies towards greater efficiency and speed.
+To develop such a parallel network, we build on UAT, integrating DUAT Wang & Li (2024b; a) to theoretically propose a foundational approach to parallel network design.
+To validate our theory, we designed a parallel network, Para-Former, based on the existing ViT architecture.
 Our contributions include:
-  • Proposing the theory of parallel deep learning networks and experimentally validating its feasibility.
-  • Analyzing the characteristics, advantages, and disadvantages of both serial and parallel networks.
-  • Investigating the issues present in deep learning from a data perspective.
-The structure of this paper is as follows: In Section 2, we first present the general form of the UAT.
-Based on the mathematical form of UAT, we define the parallel network Para-Former in Section 3.
-In Section 4, we validate the effectiveness of Para-Former in various ways: we assess its feasibility in Section 4.2, explore the impact of network depth in Section 4.3, and analyze the influence of datasets on network performances in Section 4.4.
-2 UAT
-Deep learning networks have demonstrated their powerful capabilities in numerous CV and NLP tasks.
-In the papers UAT2LLMs (Wang and Li 2024b) and UAT2CVs (Wang and Li 2024a), they have been proven that both Transformer-based networks and residual-based CNNs are specific implementations of the UAT.
-Therefore, we consider UAT to be the fundamental theory for deep learning.
-To design parallelized networks, we must start from the principles of UAT.
-Before formally designing parallel networks, we will first review the basic form of the UAT theory and then propose a network parallelization solution based on this mathematical theory.
-The UAT was initially proposed by Cybenko (Cybenko 2007), and its basic form is as follows:
+  • We propose a parallel network approach based on DUAT.
+First, we outline the direction for parallel network design based on UAT and then identify the issues associated with this design.
+Next, we present the fundamental principles and requirements for parallel network design based on DUAT.
+  • We analyze the characteristics, advantages, and disadvantages of both serial and parallel networks from the perspective of DUAT.
+  • We investigate the issues on performance present in deep learning from a data perspective and give the solution.
+The structure of this paper is as follows: In Section 2, we begin with a brief introduction to the general form of UAT (Section 2.1).
+Next, we present a method for designing parallel networks based on the UAT and discuss the prolems associated with such network designs (Section 2.2).
+To address these challenges, we introduce DUAT in Section 3, providing an initial overview in Section 3.1 before proposing the Para-former model based on DUAT in Section 3.2.
+Finally, we analyze the efficiency improvements of this parallel network in Section 3.3.
+In Section 4, we validate the effectiveness of Para-Former through various evaluations: assessing its feasibility in Section 4.2, examining the impact of network depth in Section 4.3, and analyzing the influence of the dataset on network performance in Section 4.4.
+2 UAT to Parallel Network
+2.1 Introduction of UAT
+To develop a parallel network, we start with mathematical forms of UAT.
+Before formally designing the parallelized network, we will first review the generalized mathematical forms of UAT.
+The UAT was originally introduced by Cybenko, with the following basic form:
 $G(\mathbf{x})=\sum_{j=1}^{N}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$ (1)
@@
 This indicates that when $N$ is sufficiently large, a neural network can approximate any continuous function on a closed interval.
-Hornik, Stinchcombe, and White  (Hornik, Stinchcombe, and White 1989) further proved that this function can approximate any Borel measurable function.
+Hornik et al.
+further proved that this function can approximate any Borel measurable function.
 Observing Eq.
-(1), where the function $G(\mathbf{x})$ produces scalar outputs in the real number field $\mathbb{R}$, the scenario naturally extends when $G(\mathbf{x})$ maps to $\mathbb{R}^{m}$, requiring approximation in each dimension.
+equation 1, where the function $G(\mathbf{x})$ produces scalar outputs in the real number field $\mathbb{R}$, the scenario naturally extends when $G(\mathbf{x})$ maps to $\mathbb{R}^{m}$, requiring approximation in each dimension.
 To accommodate such multi-dimensional outputs, we simply need to extend Eq.
-(1): the transformation matrix $\mathbf{W}_{j}$ is revised to the space $\mathbb{R}^{n\times m}$, the bias term $\theta_{j}$ is recast as a vector in $\mathbb{R}^{m}$, and $\alpha_{j}$ is reshaped into a matrix.
+equation 1: the transformation matrix $\mathbf{W}_{j}$ is revised to the space $\mathbb{R}^{n\times m}$, the bias term $\theta_{j}$ is recast as a vector in $\mathbb{R}^{m}$, and $\alpha_{j}$ is reshaped into a matrix.
 We call the parameters $\mathbf{W}_{j}$ and $\alpha_{j}$ in UAT as weight and $\theta_{j}$ as bias.
-Figure 1: The general description of the serial network and parallel network.
-3 UAT for Parallel Network
-3.1 Requirements for Basic Block in Parallel Network
+2.2 The Problems in Designing parallel Network Based on UAT
+Figure 1: The general description of the serial network and parallel network based on UAT.
 From Eq.
-(1), we can observe a significant characteristic: if we define different indices $j$ as the indices of layers and the corresponding layer’s mathematical form is $\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$, it is evident that the parameters between different layers do not interact with each other.
+equation 1, we can observe a significant characteristic: if we define different indices $j$ as the indices of layers and the corresponding layer’s mathematical form is $\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$, it is evident that the inputs, outputs and parameters between different layers do not interact with each other.
 Therefore, based on Eq.
-(1), we can design networks where parameters across multiple layers are independent, achieving parallel computation, like Figure 1.
+equation 1, we can design networks where parameters across multiple layers are independent, achieving parallel computation, like Figure 1.
 Figure 1.a represents a typical sequential network, which clearly requires computation to proceed in order from front to back.
@@
 If we design the parallel network solely based on Eq.
-(1), the network will face a limitation once training is complete: the function that can be approximated by the network is fixed, as the parameters in the corresponding UAT are fixed.
-So to ensure the network can dynamically approximate functions based on input, some or all of the parameters $\alpha_{j}$, $W_{j}$ and $\theta_{j}$ must be influenced by the input (referencing to UAT2LLMs and UAT2CVs), which can be simply written as $\alpha_{j}(\mathbf{x})$, $W_{j}(\mathbf{x})$ and $\theta_{j}(\mathbf{x})$.
-This allows the network to dynamically adapt and fit the function based on the input while enabling parallel computation.
-So we could design the basic block in Figure 1, like $\text{Block}_{1}\cdots\text{Block}_{n}$ are the implementation of $\alpha_{j}(\mathbf{x})\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}(\mathbf{x})\mathbf{x}+\theta_{j}(\mathbf{x})\right),i=1\cdots n$.
-Some or all the parameters $\alpha_{j}$, $W_{j}$ and $\theta_{j}$ will change according to input.
+equation 1, the network will face a limitation: Bad performance.
+This network format is a single layer perception network MinskySeymour (1969) and it is out of date because of poor performance.
+The fundamental reason is that once training is complete: the function that can be approximated by the network is fixed, as the parameters in the corresponding UAT are fixed.
+Although UAT has powerful approximation ability, the real world is complex.
+It is impossible to approximate the world using a fix function.
+3 DUAT to Parallel Network
+In Section 2.2, we have given the drawback of designing parallel networks based on UAT straightforwardly.
+So to ensure the network can dynamically approximate the world, we adopt the DUAT.
+DUAT’s general mathematical form is the same as UAT, but some or all parameters in DUAT Wang & Li (2024b; a) can dynamically change based on input.
+This allows the network to dynamically adapt and fit the function based on the input.
+3.1 Introduction of DUAT
+The DUAT was introduced in Wang & Li (2024b; a), which can be seen as an advanced form of UAT based on the residual structure.
+To introduce DUAT, we first explain the Matrix-Vector Method Wang & Li (2024b; a).
+In neural networks, there are a variety of fundamental operations, such as convolution, pooling, multi-head attention (MHA), and feed-forward networks (FFN), they are represented primarily by engineer format rather than expressed in a unified mathematical form.
 Figure 2: The process of Transformer module.
+To address this, Wang & Li (2024b; a) introduced the Matrix-Vector Method, a technique that standardizes these operations within a unified mathematical framework, as shown in Figure 2.
+In this context, $\mathbf{x}$ and $\mathbf{y}$ denote the input and output of a basic operation, with $T(\mathbf{W})$ representing various transformations, such as convolution, and $\mathbf{W}$ corresponding to the parameters of the transformation $T$.
+The transformations $T_{D}$ and $T_{P}$ are applied to convert $\mathbf{x}$, $\mathbf{y}$, and $\mathbf{W}$ into column vectors $\mathbf{x}^{\prime}$ and $\mathbf{y}^{\prime}$, and a parameter matrix $\mathbf{W}^{\prime}$, ensuring that the equation $\mathbf{W}^{\prime}\mathbf{x}^{\prime}=\mathbf{y}^{\prime}$ is satisfied.
+To clearly distinguish these transformed variables, we mark them with a prime symbol, for instance, $\mathbf{x}\mapsto\mathbf{x}^{\prime}$.
+With this standardization, we can express the operations of multilayer networks as a cohesive mathematical form, allowing for a unified representation across different network structures.
+Figure 3: The process of Transformer module.
+Following this approach, Figure 3 provides an example of the typical operations within a module in a Transformer model.
+According to the Matrix-Vector Method, MHA can be represented as Wang & Li (2024b):
+$MHA(\mathbf{x}_{i})\mapsto\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}$ (3)
+FFN can be expressed as Wang & Li (2024b):
+$FFN(\mathbf{x}_{i})\mapsto\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{b}^{\prime}_{i+1,3}\$ (4)
+According to Eq.
+equation 3 and Eq.
+equation 4, the output $\mathbf{x}^{\prime}_{1}$ of a single-layer Transformer can be expressed as Eq.
+equation 5 (see Appendix A for more details).
+For convenience, the mathematical form of Eq.
+equation 5 can be considered equivalent to the two terms $\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$ in Eq.
+equation 1.
+$\mathbf{x}^{\prime}_{1}$ $=(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,2})$ (5)
+Using Figure 3 as a reference for a typical module in a multi-layer Transformer network, we can derive that a multi-layer Transformer is effectively a DUAT function, which can be expressed as:
+$\mathbf{x}_{i+1}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,1}^{\prime})+\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$ (6)
+For the case where $i=0$, the parameters are defined as follows: $\mathbf{W}^{\prime}_{1,1}=\mathbf{W}^{\prime}_{1,1}$, $\mathbf{b}^{\prime}_{1,3}=\mathbf{b}^{\prime}_{1,3}$, $\mathbf{W}^{\prime}_{1,3}=\mathbf{W}^{\prime}_{1,3}$, $\mathbf{W}^{\prime}_{1,2}=\mathbf{W}^{\prime}_{1,2}\mathbf{W}^{\prime}_{1,1}$, and $\mathbf{b}^{\prime}_{1,2}=\mathbf{b}^{\prime}_{1,2}$.
+For cases where $i\geq 1$, we define the parameter updates as follows.
+For each $j=1,2,\ldots,i$, the updates are: $\mathbf{W}^{\prime}_{j+1,1}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{W}_{j,1}$, $\mathbf{b}^{\prime}_{j+1,3}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{b}_{j,3}+\mathbf{b}^{\prime}_{j+1,3}$, $\mathbf{W}^{\prime}_{j+1,2}=\mathbf{W}^{\prime}_{j+1,2}\mathbf{W}_{j,1}$, and $\mathbf{W}^{\prime}_{j+1,3}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{W}^{\prime}_{j,3}$.
+Additionally, for $j=2,\ldots,i+1$, the bias terms $\mathbf{b}^{\prime}_{j,2}$ are updated according to:
+$\mathbf{b}^{\prime}_{j,2}=$ $(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})$ (7)
+$+$ $\mathbf{W}^{\prime}_{j,2}\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,3}\sigma(\mathbf{W}^{\prime}_{k,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,2}).$
+Besides, parameters within the MHA mechanism also vary dynamically with the input.
+Therefore, in the DUAT associated with Transformers, the parameters $\mathbf{W}^{\prime}_{j,1}$ and $\mathbf{W}^{\prime}_{j,2}$ for $i+1\geq j\geq 1$, and $\mathbf{W}^{\prime}_{j,3}$ for $i+1>j>1$ in layer $i$, all adapt dynamically based on the input.
+3.2 Para-former
+In Section 3.1, we introduced DUAT using the Transformer model as an example.
+Following this rationale, we designed both single-depth and multi-depth parallel Transformer structures, as shown in Figures 4 and 5.
+We selected the Transformer for this design due to its robust performance and established utility across both CV and NLP domains.
 Figure 4: The structure of the Para-Former-1-n represents Para-Former with 1 depth of n layers.
 Figure 5: The structure of the Para-Former-m-n represents Para-Former with m depth of n layers.
+It is evident that Transformer-based models, whether single-layer or multi-layer, possess the capability to dynamically approximate functions based on input.
+According to Eq.
+equation 6, the parameters responsible for this dynamic approximation in a single layer are $\mathbf{W}^{\prime}_{1,1}$ and $\mathbf{W}^{\prime}_{1,2}$.
+In multi-layer structures, both weights and bias terms exhibit dynamic approximation capabilities.
+Notably, the bias terms also leverage DUAT for approximation.
+Consequently, the input-output relationship and parameters in a multi-layer Transformer structure are inherently interrelated.
+The advantage of the single-depth model lies in its reduced computation time; however, only the weight terms are dynamically changed.
+In contrast, the multi-depth model benefits from dynamic approximation in both its weights and multiple bias terms, with bias terms also employing DUAT Wang & Li (2024b).
+These approaches in networks design raises two key questions: is this network design feasible, and if so, which network performs better when the number of basic modules is identical?
+We will address these questions in Section 4.3.
+3.3 Speed-up Ratio
+We have proposed a network parallelization scheme, and in this section, we will compare the running speed of this parallelized network with that of traditional serial networks.
+It is important to note that due to differences in hardware and implementation methods, providing an objective comparison through experiments can be challenging.
+Therefore, we will focus on a theoretical analysis.
+Based on the network structure shown in Figure 4 and 5, the inference time of Para-Former is only affected by its network depth.
+Assume we use a Para-Former with a depth of $M$ and a layer count of $L$, compared to a traditional Transformer network with a depth of $N$ layers, where $L>>N$.
+In this case, the running speed of the parallel network would be $N/M$ times faster than that of the serial network.
+This is because the speed of the parallel network is solely related to the depth of the Para-Former.
+As the serial network depth increases, the acceleration effect becomes more pronounced.
+4 Experiments
+Our experimental approach is to first verify the feasibility of the parallel network theory and its adherence to the DUAT properties (The larger value of $N$, the more powerful the model becomes, resulting in better classification performance.) using several commonly used public image classification datasets (Section 4.2).
+We then explore the impact of network depth on the Para-Former performance (Section 4.3), followed by an investigation into the influence of the dataset (Section 4.4).
+We conducted experiments with various Para-Former models, where the naming convention uses the first number to indicate the depth and the second number to indicate the number of layers.
+For instance, Para-former-1-12 denotes a model with 1 depth and 12 layers.
 Cifar-10 Cifar-100 STL-10 OxfordIIITPet Flower-102
@@
 It also provides the number of categories (Classes) and the average number of images per category for each dataset (Per class).
-3.2 Para-former
-Based on the previous request on the basic module, we designed the Para-Former which is a parallel network by using the Transformer module as the basic module.
-We chose Transformer for this design because it is a well-established and widely-used network architecture, excelling in both CV and NLP.
-Another reason is that its basic mathematical form aligns with the requirements of Section 3.1.
-Before formally introducing the structure of Para-Former, let’s briefly review the Transformer structure.
-The basic structure of a Transformer is shown in Figure 2, and for convenience, we refer to it as a Transformer module.
-Typically, it is represented as follows:
-$\mathring{\mathbf{H}}=$ $\operatorname{MultiHead}(\mathbf{Q},\mathbf{K},\mathbf{V})$ (3)
-$=$ $\operatorname{Concat}\left(\hat{\mathbf{H}}_{i1},\ldots,\hat{\mathbf{H}}_{ih}\right)\mathbf{W}_{O}$
-$\hat{\mathbf{H}}_{ij}$ $=\operatorname{Attention}(\mathbf{x}_{ij}\mathbf{W}_{Qij}=\mathbf{Q}_{ij},\mathbf{x}_{ij}\mathbf{W}_{Kij}=\mathbf{K}_{ij},$ (4)
-$\mathbf{x}_{i}\mathbf{W}_{Vi}=\mathbf{V}_{ij})$
-$=\operatorname{softmax}\left(\frac{\mathbf{Q}_{ij}\mathbf{K}_{ij}^{T}}{\sqrt{M}}\right)\mathbf{V}_{ij}$
-$=\mathbf{H}_{ij}\mathbf{V}_{ij}=\mathbf{H}_{ij}[\mathbf{x}_{ij}\mathbf{W}_{Vij}]$
-where the subscript of $ij$ represents the $j$-th head of Multi-Head Attention (MHA) in the $i$-th layer.
-Please note that throughout the following text, all the subscript $i$ refer to the layer index in the network, like $\mathbf{x}_{i}$.
-According to Figure 2, assuming the input to the Feed-Forward Network (FFN) in the $i$-th layer of the Transformer, it can be expressed as:
-$FFN(\mathbf{x}_{i})=\mathbf{W}_{i,3}\sigma(\mathbf{W}_{i,2}\mathbf{x}_{i}+\mathbf{b}_{i,2})+\mathbf{b}_{i,3}\$ (5)
-The above representation mainly describes the engineering process rather than its true mathematical form.
-According to UAT2LLMs, the MHA and FFN components of Transformers can be expressed in matrix-vector form.
-We denote the matrix-vector form of variables and operations with a prime symbol (e.g., $\mathbf{x}_{i}$ becomes $\mathbf{x}^{\prime}_{i}$).
-So MHA and FFN can be written as (For more details, refer to UAT2LLMs):
-$MHA^{\prime}(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}$ (6)
-$FFN^{\prime}(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{b}^{\prime}_{i+1,3}\$ (7)
-According to Eq.
-(6) and Eq.
-(7), the output $\mathbf{x}^{\prime}_{1}$ of a single-layer Transformer can be expressed as Eq.
-(8) (see Appendix A for more details).
-The mathematical form of Eq.
-(8) can be considered equivalent to the two terms $\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$ in Eq.
+4.1 Dataset
+In pursuit of a comprehensive understanding of the efficacy and practicality of parallel models, we meticulously selected several widely-acknowledged and representative image classification datasets for an integrated comparative analysis.
+These datasets encompass a diverse array of visual content and span various levels of difficulty, thereby furnishing us with a fertile ground to probe the performance boundaries of parallel models.
+Specifically, our study centers on the following pivotal datasets: CIFAR-10, CIFAR-100 Krizhevsky (2009), STL-10 Coates et al.
+(2011), Flowers-102 Nilsback & Zisserman (2008) and Oxford-IIIT Pets Parkhi et al.
 (2012).
-Therefore, if we parallelize multiple instances of $\mathbf{x}^{\prime}_{1}$, we obtain a multi-layer parallel Transformer network, as illustrated in Figure.
-3.
-$\mathbf{x}^{\prime}_{1}$ $=(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,2})$ (8)
-Additionally, we provide the output for a two-layer Transformer, as shown in Eq.
-(9) (see Appendix A for more details).
-From Eq.
-(9), it is evident that the serial configuration of two Transformer layers also adheres to the mathematical form of UAT.
-Therefore, we can extend the Para-Former structure.
-Specifically, a block in Figure.
-1 can contain multiple serial Transformer modules (it has been proven in UAT2LLMs that the mathematical form of multi-layer Transformers is also consistent with UAT), while multiple blocks operate in parallel, as illustrated in Figure.
-4.
-The number of Transformer modules within a single block is referred to as the depth, while the total number of Transformer modules across all blocks is referred to as the network’s layer count.
-$\mathbf{x}^{\prime}_{2}$ $=(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,1})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ (9)
-$+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,2})$
-Both design strategies possess the capability to dynamically adjust model parameters based on the input.
-In Figure 3, the primary mechanism is the dynamic adjustment of MHA weight parameters based on the input.
-In contrast, Figure 4 not only leverages the dynamic adjustment of MHA weight parameters but also utilizes UAT to dynamically approximate some bias parameters.
-For instance, in Eq.
-(9), $\mathbf{b}^{\prime}_{2,2}=(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$, can be seen as a two-layer UAT.
-Refer to Appendix B to see which parameters are influenced by the input.
-So, what are advantages of these strategies?
-We will talk about these in Section 4.3.
+In Table 1, we provide more details for these datasets.
 Para-F-1-1 Para-F-1-6 Para-F-1-12 Para-F-1-24
@@
 We simply represent them as Para-F-1-1, Para-F-1-6, Para-F-1-12 and Para-F-1-24.
-3.3 Speed-up Ratio
-We have proposed a network parallelization scheme, and in this section, we will compare the running speed of this parallelized network with that of traditional serial networks.
-It is important to note that due to differences in hardware and implementation methods, providing an objective comparison through experiments can be challenging.
-Therefore, we will focus on a theoretical analysis.
-Based on the network structure shown in Figure 3 and 4, the inference time of Para-Former is only affected by its network depth.
-Assume we use a Para-Former with a depth of $M$ and a layer count of $L$, compared to a traditional Transformer network with a depth of $N$ layers, where $L>>N$.
-In this case, the running speed of the parallel network would be $N/M$ times faster than that of the serial network.
-This is because the speed of the parallel network is solely related to the depth of the Para-Former.
-As the serial network depth increases, the acceleration effect becomes more pronounced.
-4 Experiments
-Our experimental approach is to first verify the feasibility of the parallel network theory and its adherence to the UAT approximation properties (The larger the value of $N$, the more powerful the model becomes, resulting in better classification performance.) using several commonly used public image classification datasets (Section 4.2).
-We then explore the impact of network depth on the Para-Former performance (Section 4.3), followed by an investigation into the influence of the dataset (Section 4.4).
-We conducted experiments with various Para-Former models, where the naming convention uses the first number to indicate the depth and the second number to indicate the number of layers.
-For instance, Para-former-1-12 denotes a model with 1 depth and 12 layers.
-4.1 Dataset
-In pursuit of a comprehensive understanding of the efficacy and practicality of parallel models, we meticulously selected several widely-acknowledged and representative image classification datasets for an integrated comparative analysis.
-These datasets encompass a diverse array of visual content and span various levels of difficulty, thereby furnishing us with a fertile ground to probe the performance boundaries of parallel models.
-Specifically, our study centers on the following pivotal datasets: CIFAR-10, CIFAR-100 (Krizhevsky 2009), STL-10 (Coates, Ng, and Lee 2011), Flowers-102 (Nilsback and Zisserman 2008) and Oxford-IIIT Pets (Parkhi et al.
-2012).
-In Table 1, we provide more details for these datasets.
+4.2 The Feasibility of Para-Former
+To validate the feasibility and characteristics of Para-Former, we conducted training and validation on several commonly used image classification datasets, with all models trained for 500 epochs.
+The experimental design involved preprocessing inputs using the common ViT techniques of patch extraction and positional encoding.
+We then used Transformer as the basic model architecture, setting all network depths to 1 and configuring Para-Former from single-layer to multi-layer setups, as shown in Figures 4 and 5 (We evaluated the models’ performance using the accuracy metric, calculated as $\text{accuracy}=100\times\frac{\text{correct}}{\text{total}}$).
+A single-layer setup corresponds to a ViT with a depth of 1.
+The results of the models are presented in Table 2.
+It is evident from Table 2 that as the number of parallel network layers increases, the model accuracy improves progressively (Except for the result of Para-F-1-24 being worse than Para-F-1-12 on the Flower-102 dataset, this variation is within a reasonable range.).
+According to DUAT theory, increasing the network layers is equivalent to increasing $N$ in Eq.
+equation 1, thereby enhancing the approximation capability of DUAT.
+The results in Table 2 are in complete agreement with DUAT.
+Therefore, we believe that this parallel network approach is entirely feasible.
 Para-F-6-6 Para-F-2-24 Para-F-3-24 Para-F-6-24
@@
 Table 3: This table shows the predictions of Para-former-6-6, Para-former-2-24, Para-former-3-24, Para-former-6-24 on the datasets CIFAR-10, CIFAR-100, STL-10, OxfordIIITPet, and Flower-102.
-4.2 The Feasibility of Para-Former
-To validate the feasibility and characteristics of Para-Former, we conducted training and validation on several commonly used image classification datasets, with all models trained for 500 epochs.
-The experimental design involved preprocessing inputs using the common ViT techniques of patch extraction and positional encoding.
-We then used Transformer as the basic model architecture, setting all network depths to 1 and configuring Para-Former from single-layer to multi-layer setups, as shown in Figure 3 (We evaluated the models’ performance using the accuracy metric, calculated as $\text{accuracy}=100\times\frac{\text{correct}}{\text{total}}$).
-A single-layer setup corresponds to a ViT with a depth of 1.
-The results of the models are presented in Table 2.
-It is evident from Table 2 that as the number of parallel network layers increases, the model accuracy improves progressively (Except for the result of Para-F-1-24 being worse than Para-F-1-12 on the Flower-102 dataset, this variation is within a reasonable range.).
-According to UAT theory, increasing the network layers is equivalent to increasing $N$ in Eq.
-(1), thereby enhancing the approximation capability of UAT.
-The results in Table 2 are in complete agreement with UAT theory.
-Therefore, we believe that this parallel network approach is entirely feasible.
 However, the overall prediction accuracy in Table 2 is quite low. What could be the reasons for this? We believe there are two main factors affecting the model’s predictive performance:
 1.
-Model Fitting Capability: This is influenced by the model’s size and design, particularly the value of $N$ in the UAT and the degrees of freedom of the parameters.
-A larger $N$ enhances the UAT’s ability to fit functions.
-The degrees of freedom refer to the number of parameters in the UAT that are influenced by the input.
-For instance, the degrees of freedom in Equation (8) is 2, while in Equation (9), it is 6 (see Appendix B).
+Model Fitting Capability: This is influenced by the model’s size and design, particularly the value of $N$ in the DUAT and the degrees of freedom of the parameters.
+A larger $N$ enhances the DUAT’s ability to fit functions.
+The degrees of freedom refer to the number of parameters in the DUAT that are influenced by the input.
+For instance, the degrees of freedom in Eq.
+equation 5 is 2 (see Appendix B).
 The degrees of freedom represent the model’s ability to dynamically fit corresponding functions based on the input.
@@
 Additionally, the degrees of freedom vary within the Transformer: they mainly consist of weights and biases.
-The degrees of freedom for weights stem from the MHA, primarily putting linear transformations based on the input.
-The degrees of freedom for biases come from residual connections, which can dynamically fit the bias terms using UAT.
-The deeper the residual layers, the more UAT layers correspond to the bias terms, leading to stronger dynamic fitting capabilities.
+The degrees of freedom for weights stem from the MHA, which can be simply viewed as a generated parameters based the covariance matrix of the input.
+The degrees of freedom for biases come from residual connections, which can dynamically fit the bias terms using DUAT.
+The deeper the residual layers, the more DUAT layers correspond to the bias terms, leading to stronger dynamic fitting capabilities.
 As a result, a shallow serial network can achieve the fitting ability of a multi-layer parallel network with a depth of 1.
@@
 We will consider this issue in Section 4.4.
-Figure 5: OxfordIIITPet dataset’s accuracy on different classes.
-Figure 6: Flower-102 dataset’s accuracy on different classes.
 4.3 Model Depth Effect
 Based on the conclusions from Section 4.2, this section explores the impact of network depth on prediction performance.
-According to equations (8) and (9), the depth of the model primarily affects the degrees of freedom of the model parameters within the UAT.
+According to equations equation 5 and equation 6, the depth of the model primarily affects the degrees of freedom of the model parameters within the DUAT.
 Therefore, we designed different versions of the Para-Former model with 24 layers and depths of 2, 3, and 6, comparing their predictive performance against a ViT with a depth of 6.
 The results are shown in Table 3.
-The degrees of freedom for Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24 are 48, 18, 72, 72, and 72, respectively (see Appendix B).
+The degrees of freedom for Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24 are 48, 15, 72, 72, and 60, respectively (see Appendix B).
 Why does Para-F-6-6, despite having the lowest degrees of freedom, only perform worse than Para-F-6-24?
 This is because the degrees of freedom encompass both the freedom of weights and biases.
-The freedom of weights can be understood as linear transformations derived from inputs, whereas the freedom of biases results from the approximation by multiple layers of UAT, which allows for stronger dynamic fitting capabilities.
-The deeper the layers, the stronger the approximation ability.
-Thus, Para-F-6-24 performs the best, followed by Para-F-6-6, as the number of UAT layer of their bias terms is largest (see Appendix B).
+The freedom of weights can be understood as linear transformations derived from inputs covariance, whereas the freedom of biases results from the approximation by multiple layers of DUAT, which allows for stronger dynamic fitting capabilities.
+The deeper the layers, the larger number of $N$ in DUAT for bias term, and the stronger the approximation ability.
+Thus, Para-F-6-24 performs the best, followed by Para-F-6-6, as the number of DUAT layer of their bias terms is largest (see Appendix B).
 In summary, deeper models have stronger approximation capabilities, but there is a limit to how deep a model can go, necessitating a balance between network depth and layer count.
@@
 The pre-trained parameters were obtained from training the ViT model on the ImageNet dataset.
-Figure 7: The fine-tuning process from the perspective of UAT.
 After increasing the depth of the Para-Former, the results have improved to some extent.
@@
 This is due to the diversity of the data (i.e., the similarity of features within the same class) and the variability of the background.
-Figures 5 and 6 show some results from the OxfordIIITPet and Flower-102 datasets.
-Both figures indicate that for categories with diverse backgrounds and inherent feature diversity, the prediction results are poor, as shown in Figures 5.b and 6.b.
-In contrast, when the background is consistent but the object morphology is diverse, the results are relatively better, as shown in Figures 5.a and 6.a.
-The best results are achieved when both the background and object morphology are minimal, as shown in Figures 5.c and 6.c.
+Figure 6: OxfordIIITPet dataset’s accuracy on different classes.
+Figure 7: Flower-102 dataset’s accuracy on different classes.
+Figures 6 and 7 show some results from the OxfordIIITPet and Flower-102 datasets.
+Both figures indicate that for categories with diverse backgrounds and inherent feature diversity, the prediction results are poor, as shown in Figures 6.b and 7.b.
+In contrast, when the background is consistent but the object morphology is diverse, the results are relatively better, as shown in Figures 6.a and 7.a.
+The best results are achieved when both the background and object morphology are minimal, as shown in Figures 6.c and 7.c.
 Although the Flower-102 dataset has fewer images, its backgrounds are more consistent, and the differences between species are smaller.
 Conversely, animals in the OxfordIIITPet dataset can appear in various forms, whereas flowers in the Flower-102 dataset are usually photographed from relatively fixed angles, leading to better performance.
+Figure 8: The fine-tuning process from the perspective of UAT.
 Does this mean that deep learning models are not powerful enough or lack generalization and reasoning capabilities?
@@
 This view is flawed.
-Suppose the human brain operates like a UAT model; it can automatically ignore background information before seeing these images because such backgrounds are frequently encountered in daily life, and the UAT in the brain has already fitted this information.
+Suppose the human brain operates like a DUAT model; it can automatically ignore background information before seeing these images because such backgrounds are frequently encountered in daily life, and the DUAT in the brain has already fitted this information.
 Therefore, people only need to focus on specific information and then use this information to train the brain to identify similarities and differences.
-For example, a 10-year-old child’s brain, observing the world at 10 frames per second (Read and Meyer 2000), would have processed over $1.5\times 10^{9}$ high-resolution images by age 10 (Lu et al.
-2017), many containing common objects and backgrounds from daily life.
+For example, a 10-year-old child’s brain, observing the world at 10 frames per second Read & Meyer (2000), would have processed over $1.5\times 10^{9}$ high-resolution images by age 10 Lu et al.
+(2017), many containing common objects and backgrounds from daily life.
 Hence, it is unfair to criticize deep learning models for being unable to train on small datasets and for lacking universality and generalization ability.
@@
 5 Conclusion
-In this paper, we propose a parallel network based on the UAT and validate its effectiveness on large datasets.
-We analyze the characteristics of both serial and parallel networks.
-Serial networks based on residual connections, often have higher degrees of parameter freedom, leading to stronger generalization capabilities.
-However, as the network depth increases, the inference time also becomes longer.
-On the other hand, parallel networks trade computational resources for time, meaning their runtime is not affected by the number of network layers.
-This allows for the design of multi-layer parallel networks that can replace serial networks to achieve fast inference.
-Additionally, we analyze the impact and reasons data have on deep learning.
-From the UAT perspective, training on large datasets essentially enhances the model’s ability to fit diverse data.
-Appendix A The UAT Format of Transformer-based ViTs
-In this section, we will demonstrate that the mathematical form of single-layer and two-layer Transformers is equivalent to the UAT.
+In this paper, we propose design guidelines for parallel neural networks based on DUAT, which, to our knowledge, is the first work to establish principles for parallel networks.
+Based on this theory, we designed a parallel network, Para-Former, and validated its effectiveness across several commonly used datasets.
+Through DUAT, we analyzed the characteristics of both serial and parallel networks.
+Serial networks with residual connections often have higher parameter flexibility, leading to better generalization capabilities.
+However, as network depth increases, inference time also grows.
+In contrast, parallel networks trade off computational resources for time efficiency, meaning their runtime is unaffected by the number of layers.
+This allows us to design multi-layer parallel networks that can replace serial networks, achieving faster inference.
+Additionally, we examined the impact of data on deep learning performance and explored the underlying reasons.
+From the DUAT perspective, training with large datasets essentially enhances a model’s capacity to fit diverse data.
+Appendix A The DUAT Format of single layer ViT
+In this section, we will demonstrate that the DUAT form of a single-layer layer ViT .
 Firstly, the basic form of a Transformer is shown in Figure 2 of main text.
 We represent the MHA and FFN in the Transformer in matrix-vector form as Eq.
-(10) and Eq.
-(11), respectively (for specific details, refer to UAT2LLMs).
+equation 8 and Eq.
+equation 9, respectively (for specific details, refer to DUAT2LLMs).
 To distinguish between the operations and variables in the Transformer and those in the matrix-vector form, we add a prime () symbol to the operations and variables in the matrix-vector form.
 It is important to emphasize that the parameters of the MHA are influenced by the input.
-$MHA^{\prime}(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}$ (10)
-$FFN^{\prime}(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{b}^{\prime}_{i+1,3}\$ (11)
+$MHA^{\prime}(\mathbf{x}^{\prime}_{i})\mapsto\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}$ (8)
+$FFN^{\prime}(\mathbf{x}^{\prime}_{i})\mapsto\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{b}^{\prime}_{i+1,3}\$ (9)
 Next, the result of a single-layer Transformer $\mathbf{x}^{\prime}_{1}$ is:
@@
 $=\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{W}^{\prime}_{1,3}\sigma(\underline{\mathbf{W}^{\prime}_{1,2}\mathbf{W}^{\prime}_{1,1}}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})+\mathbf{b}^{\prime}_{1,3}$
-Let $\mathbf{W}^{\prime}_{1,2}=\underline{\mathbf{W}^{\prime}_{1,2}\mathbf{W}^{\prime}_{1,1}}$, so we obtain:
-$\mathbf{x}^{\prime}_{1}=(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ (13)
-where $\mathbf{W}^{\prime}_{1,1}$ and $\mathbf{W}^{\prime}_{1,2}$ are parameters influenced by the input, thus having a degree of freedom of 2.
-$\mathbf{x}^{\prime}_{2}$ $=\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{1}+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{2,2})+\mathbf{b}^{\prime}_{2,3}$ (14)
-$=(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{2,3})+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{2,2})$
-=$=\{\mathbf{W}^{\prime}_{2,1}[(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})]+\mathbf{b}^{\prime}_{2,3}\}$
-$+\mathbf{W}^{\prime}_{2,3}\sigma\{\mathbf{W}^{\prime}_{2,2}[(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})]+\mathbf{b}^{\prime}_{2,2}\}$
-=$=\{\mathbf{W}^{\prime}_{2,1}(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{b}^{\prime}_{2,3}+\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})\}$
-$+\mathbf{W}^{\prime}_{2,3}\sigma\{\mathbf{W}^{\prime}_{2,2}({\mathbf{W}}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{b}^{\prime}_{2,2}+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})\}$
-=$=\{(\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,1}}\mathbf{x}^{\prime}_{0}+\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,3}})+\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})\}$
-$+\mathbf{W}^{\prime}_{2,3}\sigma\{\underline{\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}}\mathbf{x}^{\prime}_{0}+\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}\}$
-In Eq.
-(14), we provide the matrix-vector format of 2-layer Transformer.
-Let $\mathbf{W}^{\prime}_{2,1}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,1}}$, $\mathbf{b}^{\prime}_{2,1}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,3}}$, $\mathbf{W}^{\prime}_{1,3}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}}$, $\mathbf{W}^{\prime}_{2,2}=\underline{\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}}$, and $\mathbf{b}^{\prime}_{2,2}=\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}$.
-Thus, Eq.
-(14) can be written as:
-$\mathbf{x}^{\prime}_{2}$ $=(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,1})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ (15)
-$+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,2})$
-where $\mathbf{W}^{\prime}_{2,1}$,$\mathbf{b}^{\prime}_{2,1}$,$\mathbf{W}^{\prime}_{1,2}$,$\mathbf{W}^{\prime}_{1,3}$,$\mathbf{W}^{\prime}_{2,3}$, and $\mathbf{b}^{\prime}_{2,2}$ are influenced by the input, so the degree of freedom is 6.
-Another issue is that in $\mathbf{x}^{\prime}_{1}$ and $\mathbf{x}^{\prime}_{2}$, there is a term $(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,3})$ and $(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,1})$ which doesn’t generally conform to the standard terms in UAT.
-We can consider this term as a special case and regard it as a layer in UAT.
-Firstly, when the number of network layers increases, it effectively increases $N$ in UAT.
-Given UAT’s strong approximation capabilities, adding multiple instances of this term can be viewed as a linear transformation, thus it doesn’t affect the overall approximation power of UAT.
-Furthermore, this term can be considered a special form in UAT.
-Taking ReLU as an example activation function, $(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,3})$ can be expressed as $\text{ReLU}(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,3})-\text{ReLU}(-(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,3}))$.
-In summary, we have shown that the mathematical forms of one-layer and two-layer Transformers can be seen as 2 and three-layer UAT.
 Appendix B The Degree of Freedom of Different Layers Transformer
-According to UAT2LLMs, the UAT format of multi-layer Transformer can be written as:
+According to DUAT2LLMs, the DUAT format of a multi-layer ViT can be written as:
 $\mathbf{x}_{i+1}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,1})+\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$ (11)
-where $\mathbf{b}^{\prime}_{j,2}=(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})+\mathbf{W}^{\prime}_{j,2}UAT_{j-1}^{R}$, where $UAT_{j-1}^{R}=\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,3}\sigma(\mathbf{W}^{\prime}_{k,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,2})$.
-The term $\mathbf{b}^{\prime}_{j,2}$ is approximated by a $j$-layer of UAT with $\mathbf{x}_{0}$ as input.
-All $\mathbf{W}^{\prime}_{i,1}$, $\mathbf{W}^{\prime}_{j,2}$, and $\mathbf{W}^{\prime}_{j,3}$ parameters in layer $i$, where $j=1,\ldots,i$, are dynamically adjusted based on the input.
+When $i=0$, the parameters are defined as follows: $\mathbf{W}^{\prime}_{1,1}=\mathbf{W}^{\prime}_{1,1}$, $\mathbf{b}^{\prime}_{1,3}=\mathbf{b}^{\prime}_{1,3}$, $\mathbf{W}^{\prime}_{1,3}=\mathbf{W}^{\prime}_{1,3}$, $\mathbf{W}^{\prime}_{1,2}=\mathbf{W}^{\prime}_{1,2}\mathbf{W}^{\prime}_{1,1}$, and $\mathbf{b}^{\prime}_{1,2}=\mathbf{b}^{\prime}_{1,2}$.
+For cases where $i\geq 1$, we define the parameters iteratively.
+For each $j=1,2,\ldots,i$, the updates are as follows: $\mathbf{W}^{\prime}_{j+1,1}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{W}_{j,1}$, $\mathbf{b}^{\prime}_{j+1,3}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{b}_{j,1}+\mathbf{b}^{\prime}_{j+1,3}$, $\mathbf{W}^{\prime}_{j+1,2}=\mathbf{W}^{\prime}_{j+1,2}\mathbf{W}_{j,1}$, and $\mathbf{W}^{\prime}_{j,3}=\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}^{\prime}_{i,1}\cdots\mathbf{W}^{\prime}_{j,3}$.
+Additionally, for $j=1,2,\ldots,i+1$, the bias terms $\mathbf{b}^{\prime}_{j,2}$ are updated according to:
+$\mathbf{b}^{\prime}_{j,2}=$ $(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})$ (12)
+$+$ $\mathbf{W}^{\prime}_{j,2}\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,3}\sigma(\mathbf{W}^{\prime}_{k,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,2}).$
+The term $\mathbf{b}^{\prime}_{j,2}$ is approximated by a $j-1$-layer of DUAT with $\mathbf{x}_{0}$ as input.
+The parameters $\mathbf{W}^{\prime}_{j,1}$ and $\mathbf{W}^{\prime}_{j,2}$ for $i+1\geq j\geq 1$, and $\mathbf{W}^{\prime}_{j,3}$ for $i+1>j>1$ in layer $i$, all adapt dynamically based on the input.
 Because they are affected by the weight of MHA in each layer.
-Why does the MHA operation affect the weight parameters in a linear manner?
-According to Eq.
-(12) and Eq.
-(13), this is because $\mathbf{W}^{\prime}_{1,2}=\underline{\mathbf{W}^{\prime}_{1,2}\mathbf{W}^{\prime}_{1,1}}$, where $\mathbf{W}^{\prime}_{1,1}$ represents the MHA parameters, which are influenced by the input, while $\mathbf{W}^{\prime}_{1,2}$ is a defined parameter.
-As a result, the parameter $\mathbf{W}^{\prime}_{1,2}$ in Eq.
-(13) becomes input-dependent, and this influence occurs through a linear transformation.
-Similarly, in a multi-layer Transformer, multiple layers of weight parameters are derived through similar transformations.
-Therefore, it is easy to deduce that the degrees of freedom for Transformers with 1, 2, 3, and 6 layers are 2, 6, 9, and 18, respectively.
+Therefore, it is easy to deduce that the degrees of freedom for Transformers with 1, 2, 3, and 6 layers are 2, 6, 9, and 15, respectively.
 For Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24, the degrees of freedom are shown in Table 1.
@@
 Para-F-1-24 Para-F-6-6 Para-F-2-24 Para-F-3-24 Para-F-6-24
-Degree of Freedom 2*24 18*1  6*12 9*8 18*4
+Degree of Freedom 2*24 15*1  6*12 9*8 15*4
 Table 1: The degree of freedom of Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, Para-F-6-24.
@@
 The reason lies in the nature of the degrees of freedom.
-Taking $\mathbf{W}^{\prime}_{2,2}$ and $\mathbf{b}^{\prime}_{2,2}$ from Eq.
-(15) as examples, $\mathbf{W}^{\prime}_{2,2}=\underline{\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}}$, where $\mathbf{W}^{\prime}_{1,1}$ is an input-dependent parameter from MHA, and $\mathbf{W}^{\prime}_{2,2}$ is a defined parameter, becoming fixed post-training.
-Thus, the degrees of freedom for $\mathbf{W}^{\prime}_{2,2}$ in Eq.
-(15) are imparted through a linear transformation with the MHA parameter $\mathbf{W}^{\prime}_{1,1}$.
-Additionally, $\mathbf{b}^{\prime}_{2,2}=\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}$.
-Hence, $\mathbf{b}^{\prime}_{2,2}$ in Eq.
-(15) can be understood as being approximated by a two-layer UAT, showcasing stronger dynamic transformation capabilities based on inputs.
-Due to UAT’s powerful capacity, $\mathbf{b}^{\prime}_{2,2}$ can dynamically fit specific functions based on inputs.
+Taking $\mathbf{W}^{\prime}_{2,2}$ and $\mathbf{b}^{\prime}_{2,2}$ as examples, $\mathbf{W}^{\prime}_{2,2}=\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}$, where $\mathbf{W}^{\prime}_{1,1}$ is an input-dependent parameter from MHA, and $\mathbf{W}^{\prime}_{2,2}$ is a defined parameter, becoming fixed post-training.
+Thus, the degrees of freedom for $\mathbf{W}^{\prime}_{2,2}$ is imparted through a linear transformation with the MHA parameter $\mathbf{W}^{\prime}_{1,1}$.
+Additionally, $\mathbf{b}^{\prime}_{2,2}=(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$.
+Hence, $\mathbf{b}^{\prime}_{2,2}$ can be understood as being approximated by a two-layer DUAT, showcasing stronger dynamic transformation capabilities based on inputs.
+Due to DUAT’s powerful capacity, $\mathbf{b}^{\prime}_{2,2}$ can dynamically fit specific functions based on inputs.
 In a single-layer Transformer, as seen in Eq.
-(12), $\mathbf{b}^{\prime}_{1,3}$ and $\mathbf{b}^{\prime}_{1,2}$ are defined parameters, so the number of bias terms approximated using UAT is zero.
-For a two-layer Transformer, as shown in Eq.
-(15), the number of bias terms approximated using UAT is two, specifically $\mathbf{b}^{\prime}_{2,1}$ and $\mathbf{b}^{\prime}_{2,2}$, with the UAT layers approximating these bias terms being 1 and 2, respectively.
-According to Eq.(16), in a Transformer with $i+1$ layers, $\mathbf{b}_{i+1,1}$ is approximated by a single-layer UAT, while $\mathbf{b}^{\prime}_{j,2}$ is approximated by a UAT of $j$ layers for $j>1$.
+equation 10, $\mathbf{b}^{\prime}_{1,3}$ and $\mathbf{b}^{\prime}_{1,2}$ are defined parameters, so the number of bias terms approximated using UAT is zero.
+For a two-layer Transformer, the number of bias terms approximated using DUAT is one, specifically $\mathbf{b}^{\prime}_{2,2}$ with the DUAT layers approximating this bias terms being 1.
+According to Eq.equation 11, in a Transformer with $i+1$ layers, $\mathbf{b}_{i+1,1}$ is approximated by a single-layer UAT, while $\mathbf{b}^{\prime}_{j,2}$ is approximated by a UAT of $j$ layers for $j>1$.
 It is easy to deduce that for a three-layer Transformer, the number of bias terms approximated using UAT is three, with the UAT layers being 1, 2, and 3.
-For a six-layer Transformer, the number of bias terms approximated using UAT is six, with the UAT layers being 1, 2, 3, 4, 5, and 6.
-The UAT layers of bias terms for Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24 are shown in Table 2.
+For a six-layer Transformer, the number of bias terms approximated using DUAT is six, with the DUAT layers being 1, 2, 3, 4, 5, and 6.
+The DUAT layers of bias terms for Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24 are shown in Table 2.
 Para-F-1-24 Para-F-6-6 Para-F-2-24 Para-F-3-24 Para-F-6-24
 UAT Layers (0)*24 (1+2+3+4+5+6)*1 (1+2)*12 (1+2+3)*8 (1+2+3+4+5+6)*4
-Table 2: The UAT layer count corresponding to the bias terms in Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24 is calculated by summing the UAT layers associated with each bias term within a serial Transformer.
+Table 2: The DUAT layer count corresponding to the bias terms in Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24 is calculated by summing the UAT layers associated with each bias term within a serial Transformer.
 Since each serial Transformer has multiple bias terms, with each term corresponding to different UAT layer counts, we represent the UAT layers by summing the UAT layers for each bias term.